### v3

# Dirichlet process mixtures of block $g$ priors for model selection and prediction in linear models

Abstract
This paper introduces Dirichlet process mixtures of block $g$ priors for model selection and prediction in linear models. These priors are extensions of traditional mixtures of $g$ priors that allow for differential shrinkage for various (data-selected) blocks of parameters while fully accounting for the predictors’ correlation structure, providing a bridge between the literatures on model selection and continuous shrinkage priors. We show that Dirichlet process mixtures of block $g$ priors are consistent in various senses and, in particular, that they avoid the conditional Lindley “paradox” highlighted by Som et al. (2016). Further, we develop a Markov chain Monte Carlo algorithm for posterior inference that requires only minimal ad-hoc tuning. Finally, we investigate the empirical performance of the prior in various real and simulated datasets. In the presence of a small number of very large effects, Dirichlet process mixtures of block $g$ priors lead to higher power for detecting smaller but significant effects without only a minimal increase in the number of false discoveries.
Keywords: linear model, model selection, conditional Lindley paradox, $g$ prior, continuous shrinkage prior
1 Introduction
Model selection and model averaging are foundational tasks in statistics and machine learning. Associated Bayesian procedures typically rely on the computation of Bayes factors and posterior model probabilities, whose properties are heavily dependent of the choice of priors associated with the parameters of each model. This feature makes model selection and model averaging tasks difficult in situations where standard “objective” or “default” priors (such as the reference prior, Berger et al., 2009) are improper. This is true even in well-studied settings, such as Gaussian linear models.
The literature on noninformative priors and default Bayes factors for model selection for (generalized) linear models is extensive. Examples include $g$-priors (Zellner, 1986), mixtures of $g$-priors (Zellner and Siow, 1980; Liang et al., 2008), unit information priors (Kass and Wasserman, 1995), intrinsic Bayes factors (Berger and Pericchi, 1996), non-local priors (Johnson and Rossell, 2010, 2012) and power-expected-posterior priors (Fouskakis et al., 2015; Porwal and Rodríguez, 2023), among other approaches. See Forte et al. (2018) and Consonni et al. (2018) for recent reviews.
Bayarri et al. (2012) describes a series of desiderata for default priors used for model selection and model averaging, with a particular focus on problems involving multiple linear regression. These include various forms of consistency and invariance, as well as predictive matching. More recently, Som (2014) and Som et al. (2016) suggested additional criteria related to the behavior of the Bayes factor as a subset of the significant coefficients grow to infinity. This setting is important because it serves as a proxy for situations in which effects sizes vary dramatically across covariates. Such situations arise often in practice, and they are arguably the kind of problem in which well-designed statistical methods can make a real difference. Som (2014) and Som et al. (2016) show that mixtures of $g$ priors (Liang et al., 2008) fail to satisfy their new criteria (a behavior they call the conditional Lindley paradox), and introduce mixtures of block $g$ priors, which address the issue by assigning different shrinkage parameters to preselected groups of coefficients.
In the absence of prior information to drive the a priori selection of the blocks, the methodology of Som (2014) is difficult to implement in practice. Furthermore, the prior introduced in Som (2014) assumes that the blocks of coefficients are independent a priori. When there is strong colinearity between covariates associated with “large” and “small” coefficients, the independence assumption can lead to loss of efficiency. Our first contribution in this paper is to develop Dirichlet process (DP) mixtures of block $g$ priors that allow for differential shrinkage across coefficients while fully accounting for the observed correlations among predictors and treating the blocks of covariates as an unknown parameter that must be inferred from the data. Similar approaches have been suggested in the literature at least as early as in Liang et al. (2008) but, to the best of our knowledge, they have not been pursued before, perhaps because of perceived computational challenges.
Because of our focus on differential shrinkage, the literature on continuous shrinkage priors is also relevant to our discussion. Examples of continuous shrinkage priors include the Student $t$-prior (Tipping, 2001), the Bayesian Lasso (Park and Casella, 2008; Hans, 2009), the Horseshoe prior (Carvalho et al., 2010), the Normal-Gamma prior (Griffin and Brown, 2005; Brown and Griffin, 2010), semiparametric multiple-shrinkage priors (MacLehose and Dunson, 2010), the Bayesian adaptive Lasso (Leng et al., 2014), the Dirichlet-Laplace prior (Bhattacharya et al., 2015), global-local shrinkage priors (Polson and Scott, 2012), the Beta-prime prior (Bai and Ghosh, 2018), the Horseshoe-pit prior (Denti et al., 2023), the group Inverse-Gamma Gamma prior of Boss et al. (2023), and global-local-tail priors (Lee et al., 2024) among others. Continuous shrinkage priors tend to have computational advantages, can be connected to penalized likelihood methods, and are very effective in predictive settings. However, because they place probability zero on any one value of the parameter space, variable selection can be performed only by either looking at the coverage of posterior credible intervals or by thresholding the posterior distributions of the coefficients (e.g., see Li and Pati, 2017). Both of these procedures tend to work best in settings where enough prior information is available to establish practical significance. For this reason, the literature on continuous shrinkage priors is often considered as distinct from that on priors for model selection. A second contribution of this paper is to show that DP mixtures of $g$ priors provide a unifying framework for these two strands of the literature, with canonical methods in each of the two corresponding to special cases of ours.
The remainder of the paper is organized as follows. Section 2 introduces our notation and reviews the conditional Lindley paradox. Section 3 introduces our proposed methodology and reviews its connections with the broader literature. In Section 4, we investigate the properties of the prior and the associated Bayes factors, with a particular emphasis on the criteria introduced in Bayarri et al. (2012) and Som et al. (2016). Section 5 discusses the computational implementation of our model. Section 6 and Section 7 illustrate the performance of our methodology in both simulated and real datasets. Finally, Section 8 discusses future directions for research.
2 Motivation: Bayesian variable selection and mixtures of $g$-priors
Consider a collection of linear models for the observed response vector $\boldsymbol{y}=(y_{1},\ldots,y_{n})^{T}$ based on the $n\times p$ (centered) design matrix $\boldsymbol{X}$. The collection of models is indexed by the binary vector $\boldsymbol{\gamma}=(\gamma_{1},\ldots,\gamma_{p})$, $\gamma_{j}\in\{0,1\}$, so that
$\mathcal{M}_{\boldsymbol{\gamma}}:\boldsymbol{y}=\boldsymbol{1}_{n}\beta_{0}+\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{\beta}_{\boldsymbol{\gamma}}+\boldsymbol{\epsilon},$
where $\boldsymbol{\epsilon}\sim\mathsf{N}_{n}(\boldsymbol{0},\sigma^{2}\boldsymbol{I}_{n})$, the $n$-th variate normal distribuion with mean $\boldsymbol{0}$ and covariance matrix proportional to the $n\times n$ identity matrix $\boldsymbol{I}_{n}$, $\boldsymbol{1}_{n}$ is the $n$-dimensional vector of ones, $\beta_{0}$ is an unknown intercept, $\boldsymbol{X}_{\boldsymbol{\gamma}}$ denotes the submatrix of $\boldsymbol{X}$ consisting on the columns for which $\gamma_{j}=1$, $\boldsymbol{\beta}$ is the vector of unknown regression coefficients, and $\boldsymbol{\beta}_{\boldsymbol{\gamma}}$ is the subvector of $\boldsymbol{\beta}=(\beta_{1},\ldots,\beta_{p})^{T}$ corresponding to the entries for which $\gamma_{j}=1$.
We are interested in model comparison problems among these $2^{p}$ models, as well as estimation and prediction under model uncertainty. The classical Bayesian solution to these problems involves the computation of Bayes factors of the form
$BF_{\boldsymbol{\gamma},\boldsymbol{\gamma}^{\prime}}(\boldsymbol{y})=\frac{\int f(\boldsymbol{y}\mid\beta_{0},\boldsymbol{\beta}_{\boldsymbol{\gamma}},\sigma^{2},\boldsymbol{\gamma})f(\beta_{0},\boldsymbol{\beta}_{\boldsymbol{\gamma}},\sigma^{2}\mid\boldsymbol{\gamma})\mathop{}\!\mathrm{d}\beta_{0}\mathop{}\!\mathrm{d}\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mathop{}\!\mathrm{d}\sigma^{2}}{\int f(\boldsymbol{y}\mid\beta_{0},\boldsymbol{\beta}_{\boldsymbol{\gamma}^{\prime}},\sigma^{2},\boldsymbol{\gamma}^{\prime})f(\beta_{0},\boldsymbol{\beta}_{\boldsymbol{\gamma}^{\prime}},\sigma^{2}\mid\boldsymbol{\gamma}^{\prime})\mathop{}\!\mathrm{d}\beta_{0}\mathop{}\!\mathrm{d}\boldsymbol{\beta}_{\boldsymbol{\gamma}^{\prime}}\mathop{}\!\mathrm{d}\sigma^{2}}$
for an appropriate model-specific prior $f(\beta_{0},\boldsymbol{\beta}_{\boldsymbol{\gamma}},\sigma^{2}\mid\boldsymbol{\gamma})$, which is often factorized as $f(\beta_{0},\boldsymbol{\beta}_{\boldsymbol{\gamma}},\sigma^{2}\mid\boldsymbol{\gamma})=f(\beta_{0},\sigma^{2})f(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\sigma^{2},\boldsymbol{\gamma})$. The parameters $(\beta_{0},\sigma^{2})$ are usually assigned the reference prior $f(\beta_{0},\sigma^{2})\propto\frac{1}{\sigma^{2}}$ (e.g., see Berger et al., 1998). A common choice for $f(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\sigma^{2},\boldsymbol{\gamma})$ is the so-called mixture of $g$-priors (Liang et al., 2008)
$f(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\sigma^{2},\boldsymbol{\gamma})=\int\phi\left(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\boldsymbol{0},g\sigma^{2}\left\{\boldsymbol{X}^{T}_{\boldsymbol{\gamma}}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}\right)f(g\mid\boldsymbol{\gamma})\mathop{}\!\mathrm{d}g,$
where $\phi$ denotes the density of the multivariate normal distribution and $f(g\mid\boldsymbol{\gamma})$ is a suitable hyperprior. When $f(g\mid\boldsymbol{\gamma})$ is chosen carefully (e.g., an appropriately scaled member of the Compound Confluent Hypergeometric distribution introduced in Gordy, 1998), and under mild regularity conditions, Bayes factors based on mixtures of $g$ priors have various appealing theoretical properties like model selection consistency and information consistency (Liang et al., 2008; Bayarri et al., 2012).
In spite of these strong theoretical guarantees, procedures based on mixtures of $g$ priors do suffer from some undesirable properties. For example, Som (2014) and Som et al. (2016) showed that Bayes factor based on mixtures of $g$ priors suffer from the conditional Lindley paradox. Roughly speaking, this “paradox” states that, when comparing nested models, if at least one of the regression coefficients common to both models is large relative to other coefficients present only in the bigger model, the Bayes factor will place too much weight on the smaller model irrespective of the data generating model. Consider the two models,
$\mathcal{M}_{0}:$ $\boldsymbol{y}=\boldsymbol{1}_{n}\beta_{0}+\boldsymbol{X}_{1}\boldsymbol{\beta}_{1}+\boldsymbol{\epsilon},\quad\mathcal{M}_{a}:\boldsymbol{y}=\boldsymbol{1}_{n}\beta_{0}+\boldsymbol{X}_{1}\boldsymbol{\beta}_{1}+\boldsymbol{X}_{2}\boldsymbol{\beta}_{2}+\boldsymbol{\epsilon},$ (2.1)
where $\boldsymbol{X}_{1}$ and $\boldsymbol{X}_{2}$ are $n\times p_{1}$ and $n\times p_{2}$ dimensional matrices such that $\boldsymbol{X}_{1}^{T}\boldsymbol{X}_{2}=\boldsymbol{0}$, $\boldsymbol{\beta}_{1}$ and $\boldsymbol{\beta}_{2}$ are $p_{1}$ and $p-p_{1}$ dimensional vectors, and $\boldsymbol{\epsilon}$ is the observational noise. Further, for fixed $n$, $_{1}, pp$, $\boldsymbol{X}_{1}$, $\boldsymbol{X}_{2}$, $\beta_{0}$, $\boldsymbol{\beta}_{2}\neq\boldsymbol{0}$ and $\boldsymbol{\epsilon}$, consider a sequence of vectors $\{\boldsymbol{\beta}_{1}(N):N\in\mathbb{N}\}$ and the associated sequence $\{\boldsymbol{y}(N):N\in\mathbb{N}\}$ such that $\boldsymbol{y}(N)=\boldsymbol{1}_{n}\beta_{0}+\boldsymbol{X}_{1}\boldsymbol{\beta}_{1}(N)+\boldsymbol{X}_{2}\boldsymbol{\beta}_{2}+\boldsymbol{\epsilon}$. Som et al. (2016) showed that, if $\left\|\boldsymbol{\beta}_{1}(N)\right\|\rightarrow\infty$ as $N\rightarrow\infty$, then, for the Bayes Factor for comparing $\mathcal{M}_{a}$ and $\mathcal{M}_{0}$, $BF_{a,0}(\boldsymbol{y})$, based on the hyper-$g/n$ distribution $f(g)=(1+g/n)^{-a/2}$ (Liang et al., 2008), we have $BF_{a,0}\left(\boldsymbol{y}(N)\right)\rightarrow 0$, irrespective of $\boldsymbol{X}_{1}$, $\boldsymbol{X}_{2}$, $\beta_{0}$, $\boldsymbol{\beta}_{2}$ and $\epsilon$. To illustrate the paradox, we present in Figure 1 the behavior of $\log BF_{a,0}\left(\boldsymbol{y}(N)\right)$ for 100 randomly constructed triads $(\boldsymbol{X}_{1},\boldsymbol{X}_{2},\boldsymbol{\epsilon})$ where $n=100$, $p=2$, $p_{1}=1$, $\beta_{0}=0.5,\beta_{2}=1$. We see that, in every case, $\log BF_{a,0}\left(\boldsymbol{y}(N)\right)$ seems to decrease towards -$\infty$.
Figure 1: Empirical illustration of the conditional Lindley paradox. Thin grey lines correspond to 100 simulated datasets, while the thick blue line corresponds to the average.
The conditional Lindley paradox is a consequence of the use of a common shrinkage factor $g$ and cannot be solved through alternative choices of the prior on a single shrinkage coefficient $g$. This is because, as some coefficients grow, the estimate of the common $g$ also must grow. The result is that small but non-zero coefficient end up being shrunk towards zero. To remedy this, Som (2014) propose priors that allow for different shrinkage coefficients for various blocks of parameters. In the case of $\mathcal{M}_{0}$ and $\mathcal{M}_{a}$ above, these are
$\boldsymbol{\beta}_{1}\mid g,\sigma^{2},\mathcal{M}_{0}\sim\mathsf{N}\left(\boldsymbol{0},\sigma^{2}g\left\{\boldsymbol{X}_{1}^{T}\boldsymbol{X}_{1}\right\}^{-1}\right)$ (2.2)
and
$\begin{pmatrix}\boldsymbol{\beta}_{1}\\ \boldsymbol{\beta}_{2}\end{pmatrix}\mid g_{1},g_{2},\sigma^{2},\mathcal{M}_{a}\sim\mathsf{N}\left(\begin{pmatrix}\boldsymbol{0}\\ \boldsymbol{0}\end{pmatrix},\sigma^{2}\begin{pmatrix}g_{1}\left\{\boldsymbol{X}_{1}^{T}\boldsymbol{X}_{1}\right\}^{-1}&\boldsymbol{0}\\ \boldsymbol{0}&g_{2}\left\{\boldsymbol{X}_{2}^{T}\boldsymbol{X}_{2}\right\}^{-1}\end{pmatrix}\right)$, (2.3)
where $g$, $g_{1}$ and $g_{2}$ are independent and identically distributed (i.i.d.), e.g., from a hyper-$g/n$ distribution. Som et al. (2016) showed that, under this block $g$-prior, the limit of $BF_{1,0}\left(\boldsymbol{y}(N)\right)$ as $N$ grows has a strictly positive lower bound, therefore avoiding the conditional Lindley paradox. Note, however, that implementing this strategy requires that we specify up front which groups of parameters will be assigned a common shrinkage parameter. This is problematic because the structure of the blocks can have a very big impact on the performance of the methods. Indeed, if at least one of the components of $\boldsymbol{\beta}_{2}$ goes to infinite as well, so that $0<\lim_{N\to\infty}\frac{\|\boldsymbol{\beta}_{1}\|^{2}}{\|\boldsymbol{\beta}_{2}\|^{2}}=d<\infty$ , then $\lim_{N\to\infty}BF_{a,0}(\boldsymbol{y}(N))=0$.
3 A new class of priors: Dirichlet process mixtures of block $g$ priors
In this paper we consider mixtures of priors of the form
$\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}},\sigma^{2},\boldsymbol{\gamma}\sim\mathsf{N}\left(\boldsymbol{0},\sigma^{2}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right),$ (3.1)
where $p_{\boldsymbol{\gamma}}=\sum_{j=1}^{p}\gamma_{j}$, $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}$ is a known covariance matrix that might depend on model $\boldsymbol{\gamma}$, $\boldsymbol{G}^{1/2}_{\boldsymbol{\gamma}}=\operatorname{diag}\{g_{1}^{1/2},\ldots,g_{p_{\boldsymbol{\gamma}}}^{1/2}\}$, and $g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}}$ are identically distributed. The associated marginal likelihood conditional on $\boldsymbol{\gamma}$ and $g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}}$ is given by:
$f(\boldsymbol{y}\mid\boldsymbol{\gamma},g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}})=\frac{\Gamma\left(\frac{n-1}{2}\right)}{\pi^{\frac{n-1}{2}}\sqrt{n}}\left|\boldsymbol{\Omega}_{\boldsymbol{\gamma}}\right|^{-1/2}\left[\boldsymbol{y}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{y}-n\bar{\boldsymbol{y}}^{2}\right]^{-\frac{n-1}{2}}.$ (3.2)
where $\boldsymbol{\Omega}_{\boldsymbol{\gamma}}=\boldsymbol{I}_{n}+\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}$ (see Section A of the supplementary materials). The differential-shrinkage $g$ prior can be obtained by setting $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}$. Indeed, note that the standard $g$ prior is then obtained by further setting $g_{1}=g_{2}=\cdots=g_{p_{\boldsymbol{\gamma}}}=g$.
A natural approach to modeling the $g_{j}$s is to assign them a parametric family that is flexible enough to encompass various tail behaviors. One example is
$f(g\mid\tau^{2},a,b)$ $=\frac{\Gamma(a+b+2)}{\tau^{2}\Gamma(a+1)\Gamma(b+1)}g^{b}\left(1+\frac{g}{\tau^{2}}\right)^{-a-b-2},$ $g$ $>0,$ (3.3)
where $\Gamma(z)=\int_{0}^{\infty}t^{z-1}\exp\left\{-t\right\}\mathop{}\!\mathrm{d}t$ denotes the standard Gamma function. Note that this family is defined for a,b>1a,b>-1 and $\tau^{2}>0$ and corresponds to a Beta prior on $g/(\tau^{2}+g)$. It includes, for example, the hyper-$g/n$ prior (which corresponds to $1/2\leq a\leq 0$, $b=0$ and $\tau^{2}=n$), as well as the half Cauchy distribution that underlies the Horseshoe prior (Carvalho et al., 2010) (which corresponds to a=b=1/2a=b=-1/2 and assigning $\tau^{2}$ a half-Cauchy distribution). Hence, borrowing from the literature on continuous shrinkage priors, we call priors of this type “global-local” $g$ priors, where $g_{1},\ldots,g_{p_{\gamma}}$ are “local” shrinkage parameter and $\tau^{2}$ is a “global” shrinkage parameter (which can be either known or unknown).
One potential challenge of the approach just outlined is the need to estimate what is, potentially, a very large number of different shrinkage factors. This is not only computationally costly, but the data is likely to have limited information about each of them, especially in high-correlation settings. Another challenge is that the performance of the procedure can be affected by the choice of the parametric family used to model the $g_{j}$s. In particular, recent work by Piironen et al. (2017) and Lee et al. (2024) indicates that, in the context of continuous shrinkage priors, the optimal tail behavior of $p(g_{j}\mid\tau^{2})$ might depend on the level of sparsity among coefficients.
Both of these challenges can be addressed through the use of a nonparametric specification for the distribution of the $g_{j}$s based on the Dirichlet process (Ferguson, 1973). A random distribution $H$ is said to follow a Dirichlet process prior with centering measure $H_{0}$ and concentration parameter $\alpha$, denoted $H\mid H_{0},\alpha\sim\mathsf{DP}(\alpha,H_{0})$, if it admits a representation of the form
$H(\cdot)=\sum_{k=1}^{\infty}w_{k}\delta_{g^{*}_{k}}(\cdot),$ (3.4)
where $\delta_{a}$ denotes a point mass at $a$, $g^{*}_{1},g^{*}_{2},\ldots$ is an i.i.d. sequence with $g^{*}_{k}\sim H_{0}$, and $w_{k}=v_{k}\prod_{l<k}(1-v_{l})$, with $v_{1},v_{2},\ldots$ another i.i.d. sequence with $v_{k}\sim\mathsf{beta}(1,\alpha)$ (Sethuraman, 1994). Because the samples from a Dirichlet process are almost surely discrete distributions, if $g_{1},\ldots,g_{c}$ is an i.i.d. sample from a random $H\mid H_{0},\alpha\sim\mathsf{DP}(\alpha,H_{0})$, there is a positive probability of ties among the $g_{i}$s. In fact, their joint distribution can be described through two sequences, $\tilde{g}_{1},\tilde{g}_{2},\ldots$ i.i.d. such that $\tilde{g}_{k}\sim H_{0}$, and $\xi_{1},\ldots,\xi_{c}$ such that $\xi_{1}=1$ and
$\xi_{j}\mid\xi_{j-1},\ldots,\xi_{1},\alpha$ $\sim\sum_{k=1}^{K^{j-1}}\frac{m_{k}^{j}}{j+\alpha-1}\delta_{k}+\frac{\alpha}{j+\alpha-1}\delta_{K^{j-1}+1},$ $j\geq 2,$
where $K^{j-1}=\max_{j^{\prime}<j}\{\xi_{j^{\prime}}\}$ and $m_{k}^{j}=\sum_{j^{\prime}<j}\mathbb{I}(\xi_{j^{\prime}}=k)$ (Blackwell and MacQueen, 1973). The value of $g_{j}$ can then be recovered from those of $\tilde{g}_{1},\tilde{g}_{2},\ldots$ and $\xi_{1},\ldots,\xi_{c}$ through the relationship $g_{j}=\tilde{g}_{\xi_{j}}$. The vector $\boldsymbol{\xi}$ defines a partition $\rho=\{S_{1},\ldots,S_{K}\}$ of the set $\mathcal{I}=\{1,\ldots,c\}$ such that $\cup_{k=1}^{K}S_{k}=\mathcal{I}$, $S_{k}\cap S_{k^{\prime}}=\emptyset$ for $k\neq k^{\prime}$, and $|S_{k}|=m_{k}$ is the number of elements in $S_{k}$, so that $i\in S_{k}$ if and only if $\xi_{i}=k$ and $f(\rho\mid\alpha)=\frac{\Gamma(\alpha)}{\Gamma(\alpha+c)}\alpha^{K}\prod_{k=1}^{K}\Gamma(m_{k})$.
In our setting, we let $g_{j}\mid H$ i.i.d. for $j=1,\ldots,p_{\boldsymbol{\gamma}}$ and $H\mid\alpha\sim\mathsf{DP}(\alpha,H_{0})$, where $H_{0}$ is the distribution associated with (3.3), and $\alpha$ is assigned the parameterization-invariant prior introduced in Rodríguez (2013), which has density $f(\alpha\mid\boldsymbol{\gamma})=\sqrt{\frac{1}{\alpha}\sum_{j=1}^{p_{\boldsymbol{\gamma}}-1}\frac{j}{(\alpha+j)^{2}}}$. The use of this prior circumvents the need to elicit prior information about the number of groups into which the coefficients are expected to be clustered. The resulting prior,
$p(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid a,b,\tau^{2},\boldsymbol{\gamma})=\int\phi\left(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\boldsymbol{0},\sigma^{2}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right)\\ f\left(\tilde{\boldsymbol{g}}\mid\rho,\gamma,\tau^{2}=n,a,b\right)f(\rho\mid\gamma,\alpha)f(\alpha\mid\boldsymbol{\gamma})\mathop{}\!\mathrm{d}\tilde{\boldsymbol{g}}\mathop{}\!\mathrm{d}\rho\mathop{}\!\mathrm{d}\alpha,$ (3.5)
is a Dirichlet process mixture of block $g$ priors.
Because there might be ties among the $g_{j}$s, the model implicitly defines a partition of the coefficients in which those assigned to the same group share a common shrinkage factor. The concentration parameter $\alpha$ controls the prior distribution on the partitions, with $\alpha\to 0$ leading to the standard (mixture of) $g$ priors, and $\alpha\to\infty$ leading back to global-local $g$ priors defined above where each coefficient is assigned its own shrinkage factor. Since the model treats both $\alpha$ and $\rho$ as unknown, the model is able to learn an appropriate partition of the coefficients as it performs model selection. Furthermore, the use of a nonparametric prior for $H$ implies that the model is potentially capable of learning from the data the shape of the distribution of the shrinkage factors,which can alleviate concerns about the specific choice of the hyperparameters $a$, $b$ and $\tau^{2}$.
3.1 Unifying continuous shrinkage and variable selection priors
Traditionally, frameworks based on continuous shrinkage priors have recognized that differential shrinkage might be needed to attain optimal performance, but until recently (e.g., Boss et al., 2023) they have tended to downplay the need to account for co-linearity among covariates. On the other hand, the literature on priors for model selection has, from the very beginning, acknowledged the need to account for colinearity, but has been slower to recognize the need for differential shrinkage, perhaps because of the computational challenges involved. Our approach provide a unifying framework for thinking about these two strands of the literature. To see this, consider a slightly less general spike-and-slab version of our prior where $\boldsymbol{\beta}\mid\sigma^{2},\tilde{g}_{1},\ldots,\tilde{g}_{p},\boldsymbol{\gamma}\sim\mathsf{N}\left(\boldsymbol{0},\sigma^{2}\tilde{\boldsymbol{G}}^{1/2}\left\{\boldsymbol{\Gamma}\boldsymbol{\Sigma}^{-1}\boldsymbol{\Gamma}\right\}^{-}\tilde{\boldsymbol{G}}^{1/2}\right)$, $\boldsymbol{\Gamma}=\operatorname{diag}\{\gamma_{1},\ldots,\gamma_{p}\}$, $\tilde{\boldsymbol{G}}=\operatorname{diag}\{\tilde{g}_{\xi_{1}},\ldots,\tilde{g}_{\xi_{p}}\}$, and $A^{-}$ represents the Moore-Penrose inverse of $A$. Different choices of $\boldsymbol{\Sigma}$ and of priors on $\boldsymbol{\gamma}$ and $\rho$ lead to various well-known procedures. For example, as we noted before, the DP mixture of block $g$ priors includes the standard $g$ prior and the “global-local” block $g$ prior as special (limit) cases. Furthermore, when $\boldsymbol{X}$ is orthogonal and the grouping variable $\rho$ is treated as known, it also includes the block $g$ prior of Som (2014) as a special case.
On the other hand, if we fix the model to $\boldsymbol{\gamma}=(1,1,\ldots,1)$, either $\boldsymbol{X}$ is orthonormal or we take $\boldsymbol{\Sigma}=\boldsymbol{I}$, and $\tau^{2}$ is given a hyperprior, the DP mixture of block $g$ priors corresponds to the Horseshoe Pit mixture prior of Denti et al. (2023). As a consequence, it also includes a number of traditional continuous shrinkage priors such as those in Park and Casella (2008), Brown and Griffin (2010), Carvalho et al. (2010), Leng et al. (2014), Bhattacharya et al. (2015) and Bai and Ghosh (2018). Also, under the full model and block orthogonality and a known group structure, our framework also includes the Group Inverse Gamma Gamma shrinkage prior of Boss et al. (2023).
4 Properties of DP mixtures of block $g$ prior
4.1 Tail behavior
Since, marginally, samples from a distribution generated by a Dirichlet process follow the baseline measure (e.g., see Blackwell and MacQueen, 1973 or Antoniak, 1974), the marginal distribution for the $l$-th entry of $\boldsymbol{\beta}_{\boldsymbol{\gamma}}$ under a DP mixture of block $g$ priors is given by $f(\boldsymbol{\beta}_{\boldsymbol{\gamma},l}\mid\tau^{2},a,b,\gamma)=\int\mathsf{N}(\boldsymbol{\beta}_{\boldsymbol{\gamma},l}\mid 0,g\kappa_{\gamma,l,l}\sigma^{2})f(g\mid\tau^{2},a,b)\mathop{}\!\mathrm{d}g$, where $\kappa_{\gamma,l,l}$ is the $l$-th diagonal entry of $\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}$. The tail behavior of this type of marginal distributions was studied in Boss et al. (2023) (see their Theorem 2.1). In particular, the index of regular variation of the marginal prior is $\omega=-2b-3$, i.e.,
$\lim_{\boldsymbol{\beta}_{\boldsymbol{\gamma},l}\to\infty}\frac{f(t\boldsymbol{\beta}_{\boldsymbol{\gamma},l}\mid\tau^{2},a,b,\gamma)}{f(\boldsymbol{\beta}_{\boldsymbol{\gamma},l}\mid\tau^{2},a,b,\gamma)}=t^{-2b-3}.$
This implies that the our prior has heavy (polynomial) tails and point estimators derived from our procedures is robust, in the sense of having bounded influence in the case of likelihood-prior conflict. However, note that the contour plots associated with the DP mixture of block $g$ priors are not elliptical, unlike those of the standard $g$ prior (please see Section B of the supplementary materials). These non-elliptical contours enable non-uniform shrinkage by the DP block $g$ priors, which in turn is what enables them to skirt the conditional Lindley paradox.
4.2 Information consistency of Bayes factors
For fixed $n$, $p$ and $\boldsymbol{X}$ that is full rank, consider a sequence of observations $\boldsymbol{y}(1),\boldsymbol{y}(2),\ldots$ such that $\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}(N)\|\to\infty$ and $N\to\infty$, where $\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}(N)=\left[\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right]^{-1}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{y}(N)$ is the maximum likelihood estimator or $\boldsymbol{\beta}_{\boldsymbol{\gamma}}$ based on $\boldsymbol{y}(N)$. The Bayes factor $BF_{\boldsymbol{\gamma},\boldsymbol{0}}$ is information consistent if $BF_{\boldsymbol{\gamma},\boldsymbol{0}}\left(\boldsymbol{y}(N)\right)\to\infty$ as $N\to\infty$.
Bayes factors under standard mixtures of $g$ priors are known to be information consistent under appropriate conditions on the prior on $g$. The following theorem, which is analogous to Theorem 2 in Liang et al. (2008), establishes general conditions on the joint prior on $g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}}$ that ensure information consistency for general mixtures of block $g$ priors.
Theorem 4.1.
Let $\nu_{+}$ be the largest eigenvalue of $\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}$ and $\lambda_{-}(\boldsymbol{G}_{\boldsymbol{\gamma}})$ be the smallest eigenvalue of $\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}-\left[\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\}^{-1}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right]^{-1}$. The mixing prior $f(g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}})$ leads to Bayes factors that are information consistent if
$\int\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right|^{-1/2}\\ \left[1-\frac{\lambda_{-}(\boldsymbol{G}_{\boldsymbol{\gamma}})}{\nu_{+}}\right]^{-\frac{n-1}{2}}f(g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}})\mathop{}\!\mathrm{d}g_{1}\ldots\mathop{}\!\mathrm{d}g_{p_{\boldsymbol{\gamma}}}=\infty.$ (4.1)
for all $p_{\boldsymbol{\gamma}}\leq p$.
The proof is included in Section C of the supplementary materials. A slightly simpler condition that applies to DP mixtures of block $g$ priors is the following:
Theorem 4.2.
A sufficient condition for the DP mixture of block $g$ priors to lead to Bayes factors that are information consistent is for the density of the centering measure, $f(g\mid\tau^{2},a,b)$, to satisfy $\int(1+g_{j})^{(n-1-p_{\boldsymbol{\gamma}})/2}f(g_{j}\mid\tau^{2},a,b)\mathop{}\!\mathrm{d}g_{j}=\infty$ for all $p_{\boldsymbol{\gamma}}\leq p$.
See Section C of the supplementary materials. This is the same condition in Theorem 2 of Liang et al. (2008). Hence, this result just indicates that any mixing distribution for $g$ that leads to information-consistent Bayes factors under a standard mixture of $g$ priors also leads to information consistent Bayes factors under the DP mixture of block $g$ priors.
4.3 Information consistency of block structures
One of the key motivations to consider DP mixtures of block $g$ priors is the desire to avoid having to decide a priori on an appropriate partition for the covariates. In this Section, we show that, when the design matrix $\boldsymbol{X}$ is orthogonal and the true coefficients have very different sizes, our prior assigns coefficients of different sizes to separate clusters with high probability. Before proceeding with our main result, we need to introduce the concept of refinement of a partition (sometimes called a fragmentation, e.g., see Bertoin, 2006).
Definition 4.1 (Refinement of a partition).
Let $\rho=\{S_{1},\dots,S_{K}\}$ and $\rho^{\prime}=\{S_{1}^{\prime},\dots,S_{K^{\prime}}^{{}^{\prime}}\}$ denote two partition of a set $\mathcal{I}=\{1,\ldots,c\}$ with $K$ and $K^{\prime}$ unique blocks respectively, such that $1\leq K^{\prime}\leq K\leq c$. Then, $\rho$ is said to be a refinement of $\rho^{\prime}$, denoted by $\rho\prec\rho^{\prime}$, if and only if for every $S_{k}\in\rho$ there exist a $S_{j}^{{}^{\prime}}\in\rho^{\prime}$ such that $S_{k}\subseteq S_{j}^{{}^{\prime}}$.
Theorem 4.3.
Let $\boldsymbol{X}$ be a full rank, centered, orthogonal design matrix of size $n\times p$, and $\boldsymbol{X}_{1}$ and $\boldsymbol{X}_{2}$ be two non-overlapping submatrices of sizes $n\times p_{1}$ and $n\times p_{2}$ with $p_{1}>0$, $p_{2}>0$ and $p_{1}+p_{2}\leq p$. Denote by $\mathcal{I}_{1}=\{j^{(1)}_{1},\ldots,j^{(1)}_{p_{1}}\}$ the set of indexes associated with the columns of $\boldsymbol{X}$ included in $\boldsymbol{X}_{1}$ and $\mathcal{I}_{2}=\{j^{(2)}_{1},\ldots,j^{(2)}_{p_{2}}\}$ the columns associated with $\boldsymbol{X}_{2}$, so that $\mathcal{I}=\mathcal{I}_{1}\cup\mathcal{I}_{2}$ and $\mathcal{I}_{1}\cap\mathcal{I}_{2}=\emptyset$. Consider now an asymptotic regime where, for fixed $n$, $p_{1}$, $p_{2}$, $\boldsymbol{X}_{1}$, $\boldsymbol{X}_{2}$, $\beta_{0}$, $\boldsymbol{\beta}_{2}$ and $\boldsymbol{\epsilon}$, a sequence of observations $\{\boldsymbol{y}(N):N\in\mathbb{N}\}$ is generated as $\boldsymbol{y}(N)=\boldsymbol{1}_{n}\beta_{0}+\boldsymbol{X}_{1}\boldsymbol{\beta}_{1}(N)+\boldsymbol{X}_{2}\boldsymbol{\beta}_{2}+\boldsymbol{\epsilon}$, where $\{\boldsymbol{\beta}_{1}(N):N\in\mathbb{N}\}$ is a sequence such that, $\beta_{j}^{2}(N)\sim\mathcal{O}(N)$ for all $j\in\mathcal{I}_{1}$. If $\rho_{0}=\{\mathcal{I}_{1},\mathcal{I}_{2}\}$, then
  1. 1.
    For $\rho\nprec\rho_{0}$, $\lim_{N\to\infty}\frac{f\left(\boldsymbol{y}(N)\mid\rho\right)}{f\left(\boldsymbol{y}(N)\mid\rho_{0}\right)}=0$, and
  2. 2.
    For $\rho\prec\rho_{0}$, $\lim_{N\to\infty}\frac{f\left(\boldsymbol{y}(N)\mid\rho\right)}{f\left(\boldsymbol{y}(N)\mid\rho_{0}\right)}=c_{\rho}$, with $0<c_{\rho}<\infty$.
The proof can be seen in Section D of the supplementary materials.
4.4 Conditional Lindley paradox
In addition to being important on its own right, the previous result allows us to show that, when the design matrix $\boldsymbol{X}$ is orthogonal, the Bayes factors based on DP mixtures of block $g$ priors avoid the conditional Lindley paradox.
Theorem 4.4.
Let $\boldsymbol{X}$ be a full rank, centered, orthogonal design matrix of size $n\times p$, and $\boldsymbol{X}_{1}$ and $\boldsymbol{X}_{2}$ be two non-overlapping submatrices of sizes $n\times p_{1}$ and $n\times p_{2}$ with $p_{1}>0$, $p_{2}>0$ and $p_{1}+p_{2}=p$, and consider the pair of models $\mathcal{M}_{\boldsymbol{\gamma}_{0}}:\boldsymbol{y}=\boldsymbol{1}_{n}\beta_{0}+\boldsymbol{X}_{1}\boldsymbol{\beta}_{1}+\boldsymbol{\epsilon}$ and $\mathcal{M}_{\boldsymbol{\gamma}_{a}}:\boldsymbol{y}=\boldsymbol{1}_{n}\beta_{0}+\boldsymbol{X}_{1}\boldsymbol{\beta}_{1}+\boldsymbol{X}_{2}\boldsymbol{\beta}_{2}+\boldsymbol{\epsilon}$. Denote by $\mathcal{I}_{1}=\{j^{(1)}_{1},\ldots,j^{(1)}_{p_{1}}\}$ the set of indexes associated with the columns of $\boldsymbol{X}$ included in $\boldsymbol{X}_{1}$ and $\mathcal{I}_{2}=\{j^{(2)}_{1},\ldots,j^{(2)}_{p_{2}}\}$ the columns associated with $\boldsymbol{X}_{2}$, so that $\mathcal{I}=\mathcal{I}_{1}\cup\mathcal{I}_{2}$ and $\mathcal{I}_{1}\cap\mathcal{I}_{2}=\emptyset$.
For fixed $n$, $p_{1}$, $p_{2}$, $\boldsymbol{X}_{1}$, $\boldsymbol{X}_{2}$, $\beta_{0}$, $\boldsymbol{\beta}_{2}$ and $\boldsymbol{\epsilon}$, let $\{\boldsymbol{\beta}_{1}(N):N\in\mathbb{N}\}$ be a sequence such that $\beta_{j}^{2}(N)\sim\mathcal{O}(N)$ as $N\to\infty$ for all $j\in\mathcal{I}_{1}$, and $\{\boldsymbol{y}(N):N\in\mathbb{N}\}$ be the associated sequence generated by setting $\boldsymbol{y}(N)=\boldsymbol{1}_{n}\beta_{0}+\boldsymbol{X}_{1}\boldsymbol{\beta}_{1}(N)+\boldsymbol{X}_{2}\boldsymbol{\beta}_{2}+\boldsymbol{\epsilon}$. Then, for the Bayes factor based on a DP mixture of block $g$ priors under the hyper-$g/n$ mixture distribution we have
$\lim_{N\to\infty}BF_{\boldsymbol{\gamma}_{a},\boldsymbol{\gamma}_{0}}\left(\boldsymbol{y}(N)\right)>0$
for any any $\{\boldsymbol{y}(N):N\in\mathbb{N}\}$ and any pair of models $\boldsymbol{\gamma}_{0}$ and $\boldsymbol{\gamma}_{a}$.
The proof, which can be found in Section E of the supplementary materials, only requires that the prior on the partition $\rho$ puts positive probability on at least one refinement of the true partition $\rho_{0}=\{\mathcal{I}_{1},\mathcal{I}_{2}\}$. This is trivially true for the Dirichlet process, whether its concentration parameter is fixed or treated as unknown and given a prior distribution. In fact, this is a very mild requirement that is also satisfied by the Bayes factor constructed under the global-local $g$ prior, and one that does not require that the partition model be able to identify the true partition $\rho_{0}$. On the other hand, orthogonality plays a key role in the proof of this theorem, just like block orthogonality is key to similar proofs in Som (2014). Simulation studies suggest that the result hold in the non-orthogonal case.
4.5 Model selection consistency
Model selection consistency refers to the ability of the procedure to choose the correct model as the sample size grows. DP mixtures of block $g$ priors are model selection consistent in the fixed $p$ regime (see Section F of the supplementary materials).
Theorem 4.5.
Assume that a sequence of observations $y_{1},y_{2},\ldots$ is generated from some model $\boldsymbol{\gamma}_{T}\in\{0,1\}^{p}$ (i.e., one of the models considered by our procedure), and that $p$ is fixed. Also, assume the following regularity conditions:
  • (i)
    the column space $\mathcal{C}(\boldsymbol{X})$ does not contain $\boldsymbol{1}_{n}$.
  • (ii)
    The sequence of covariate vectors $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots$ are such that $\|\boldsymbol{x}_{i}\|^{2}$ is bounded by a constant for all $i=1,2,\ldots$.
  • (iii)
    The smallest eigenvalue of $\boldsymbol{X}^{T}\boldsymbol{X}/n$ is lower bounded by a positive constant for all $n$.
Then, under the DP mixture block $g$ priors with $\tau^{2}=n$, $\lim_{n\to\infty}\mathsf{Pr}(\boldsymbol{\gamma}=\boldsymbol{\gamma}_{T}\mid\boldsymbol{y})=1$, as long as the prior on models satisfies $f(\boldsymbol{\gamma}_{T})>0$.
4.6 Intrinsic consistency
Under slightly more stringent regularity conditions than those required for model selection consistency, DP mixtures of block $g$ priors are also intrinsically consistent.
Theorem 4.6.
Assume that, as $n$ grows, the columns $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots$ of the design matrix satisfy either of the following two conditions for a finite, positive definite matrix $\boldsymbol{\Lambda}$:
  • (i)
    If $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots$ forms a deterministic sequence, then $\frac{1}{n}\boldsymbol{X}^{T}\boldsymbol{X}\underset{n\to\infty}{\longrightarrow}\boldsymbol{\Lambda}$.
  • (ii)
    If $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots$ are random, then they are independent and identically distributed from a distribution with mean $\boldsymbol{0}$ and covariance $\boldsymbol{\Lambda}$.
Then, the DP mixture of block $g$ priors with $\tau^{2}=n$ converges to a proper, non-degenerate intrinsic prior of the form
$f(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid a,b,\boldsymbol{\gamma})=\int\phi\left(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\boldsymbol{0},\sigma^{2}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{\Lambda}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right)\\ f\left(\tilde{\boldsymbol{g}}\mid\rho,\gamma,\tau^{2}=n,a,b\right)f(\rho\mid\boldsymbol{\gamma},\alpha)f(\alpha\mid\boldsymbol{\gamma})\mathop{}\!\mathrm{d}\tilde{g}_{1}\ldots\mathop{}\!\mathrm{d}\tilde{g}_{p_{\boldsymbol{\gamma}}}\mathop{}\!\mathrm{d}\rho\mathop{}\!\mathrm{d}\alpha.$
5 Computation
It is possible to construct MCMC algorithms for model selection under the DP mixtures of block $g$ prior that require very minimal tuning. To do so, we take advantage of the conditional conjugacy of the priors and, when possible, we integrate out the intercept $\beta_{0}$, the vector of regression coefficients $\boldsymbol{\beta}_{\boldsymbol{\gamma}}$ and/or the variance $\sigma^{2}$ when deriving conditional posteriors. Additionally, we represent the shrinkage coefficients $g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}}$ in terms of their unique values $\tilde{\boldsymbol{g}}_{\boldsymbol{\gamma}}=(\tilde{g}_{1},\ldots,\tilde{g}_{K_{\boldsymbol{\gamma}}})$ and the group indicators $\boldsymbol{\xi}_{\boldsymbol{\gamma}}=(\xi_{1},\ldots,\xi_{p_{\boldsymbol{\gamma}}})$ (recall Section 3). The resulting posterior takes the form
$f(\boldsymbol{\gamma},\tilde{\boldsymbol{g}},\boldsymbol{\xi},\alpha\mid\boldsymbol{y})\propto f(\boldsymbol{y}\mid\boldsymbol{\gamma},\tilde{\boldsymbol{g}},\boldsymbol{\xi})f\left(\tilde{\boldsymbol{g}}\mid\boldsymbol{\gamma}\right)f(\boldsymbol{\xi}\mid\boldsymbol{\gamma},\alpha)f(\alpha|\boldsymbol{\gamma})f(\boldsymbol{\gamma}),$
where $f(\boldsymbol{y}\mid\boldsymbol{\gamma},\tilde{g}_{1},\ldots,\tilde{g}_{K_{\boldsymbol{\gamma}}},\xi_{1},\ldots,\xi_{p_{\boldsymbol{\gamma}}})$ corresponds to (3.2) with $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}$, and $p(\boldsymbol{\gamma})$ is an appropriate prior on the space of models, e.g., a Beta-Binomial prior $f(\boldsymbol{\gamma})=\frac{\Gamma(c+d)}{\Gamma(c)\Gamma(d)}\frac{\Gamma(c+p_{\boldsymbol{\gamma}})\Gamma(d+p-p_{\boldsymbol{\gamma}})}{\Gamma(c+d+p)}$.
[…]
6 Simulation studies
6.1 Conditional Lindley paradox
Our first simulation study replicates the setting used to construct Figure 1 and shows empirical evidence supporting the theoretical results discussed in Sections 4.3 and 4.4. We consider a total 150 simulations, each of which involves a sequence of datasets generated under model $\mathcal{M}_{a}:y_{i}=\beta_{0}+\beta_{1}x_{i,1}+\beta_{2}x_{i,2}+\epsilon_{i}$ for $i=1,\ldots,100$. All elements of a given sequence of datasets share the same values of $\beta_{0}$, $\beta_{1}$, $\boldsymbol{x}_{1}=(x_{1,1},\ldots,x_{100,1})^{\prime}$, $\boldsymbol{x}_{2}=(x_{1,2},\ldots,x_{100,2})^{\prime}$ and $\boldsymbol{\epsilon}=(\epsilon_{1},\ldots,\epsilon_{100})^{\prime}$. In particular, we set $\beta_{0}=0.5$, $\beta_{1}=1$ and generate $\epsilon$ from a standard multivariate Gaussian distribution and each of the pairs $(x_{i,1},x_{i,2})^{\prime}$ from a zero-mean bivariate Gaussian distribution with unit marginal standard deviations and correlation $\eta$. The datasets within each sequence are then constructed by considering a grid of values for $\beta_{2}$ in the interval $[0,240]$. We are interested in the behavior of $B_{a,0}$, the Bayes factor under the DP mixture of $g$ priors that uses the hyper-$g/n$ as the baseline measure comparing the true model $\mathcal{M}_{a}$ against the simpler model $\mathcal{M}_{0}:y_{i}=\beta_{0}+\beta_{1}x_{i,1}+\epsilon_{i}$. We also investigate the behavior of $\mathsf{Pr}(\xi_{1}\neq\xi_{2}\mid\boldsymbol{y})$, the posterior probability that the model assigns different shrinkage parameters to each variable in the model (recall Section 4.3).
Figure 2 shows the results of this simulation study for $\eta=0$ and $\eta=0.5$. Compared to Figure 1, the curves are somewhat noisy. This is an artifact of the Monte Carlo noise introduced by our MCMC algorithm (the Bayes factor depicted in 1 is available in “closed form”). With that caveat in mind, we note that the curves for $\log\left(B_{a,0}(\boldsymbol{y})\right)$ decrease as $\beta_{2}$ increases but, unlike Figure 1, both seem to stabilize towards an asymptote. This agrees with the behavior predicted by Theorem 4.4. Similarly, and as predicted by Theorem 4.3, we can see that $\mathsf{Pr}(\xi_{1}\neq\xi_{2}\mid\boldsymbol{y})$ seems to converge to 1 as $\beta_{2}$ grows.
(a) Log Bayes factor, $\eta=0$
(b) $\mathsf{Pr}(\xi_{1}\neq\xi_{2}\mid\boldsymbol{y})$,, $\eta=0$
(c) Log Bayes factor, $\eta=0.5$
(d) $\mathsf{Pr}(\xi_{1}\neq\xi_{2}\mid\boldsymbol{y})$, $\eta=0.5$.
Figure 2: Behavior of $\log\left(B_{a,0}(\boldsymbol{y})\right)$ (left column) and $\mathsf{Pr}(\xi_{1}\neq\xi_{2}\mid\boldsymbol{y})$ (right column) under the DP mixture of block $g$ priors in our first simulation study. Each thin grey line corresponds to one replicate of the simulation, while the thicker blue line corresponds to the mean curve. Figures in the top row correspond to design matrices generated under $\eta=0$, while the bottom row corresponds to $\eta=0.5$
6.2 Model selection, estimation and prediction performance
We conducted a second simulation study to compare the model selection, estimation, and prediction performance of procedures based on DP mixtures of block $g$ priors with that of competing procedures. We considered three versions of DP mixtures of block $g$ priors: one using a hyper-$g$ base measure (labeled “DP block-$g$ ($\tau^{2}=1$)”), a “unit information” version using a hyper-$g/n$ base measure (labeled “DP block-$g$ ($\tau^{2}=n$)”), and one using a scaled hyper-$g$ base measure where the scale is in turn given a half-Cauchy hyperprior (labeled “DP block-$g$ ($\tau^{2}\sim\mathsf{HC}$) ”). This second study, the setup of which is inspired by Denti et al. (2023), assumes $n=500$ for all datasets and, as before, the vectors of covariates associated with the each observation are generated from a zero-mean multivariate normal distribution with unit marginal variances and correlation $\eta$ across all pairs of covariates. We consider six scenarios that arise from combining three different values for the total number of covariates ($p=250$, $p=500$ and $p=750$) and two different levels of multicolinearity ($\eta=0$ and $\eta=0.9$). For all three values of $p$, 100100 of the coefficients are randomly sampled from a normal distribution with mean 0 and standard deviation 10 (we consider these “large” coefficients), $100$ are randomly sampled from a standard normal distribution (the “small” coefficients), and the remainder are set to 0 (the null coefficients). For each of the six scenarios, we generate 100 datasets.
In terms of competing approaches, we consider the following: (a) a standard hyper-$g/n$ mixture of $g$ priors (which we label “g-prior” in the sequel); (b) a version of the block $g$ prior of Som (2014) with a hyper-$g/n$ hyperprior and known blocking structures where the covariates are allocated to $K=2$ groups: one made of all the covariates associated with large coefficients plus half, randomly chosen variables associated with null coefficients, and another one made of the rest (labeled “Som et al. ($\tau^{2}=n$, K=2K=2)”); (c) a version of the block $g$ prior of Som (2014) with $K=3$ fixed groups of covariates: one made of all the covariates associated with large coefficients, one made of the covariates associated with small coefficients, and one made of the covariates associated with null coefficients (labeled “Som et al. ($\tau^{2}=n$, K=3K=3)”); (d) three versions of “global-local” $g$-priors with distinct but identically distributed shrinkage parameters for each coefficient, one using a hyper-$g$ prior (labeled “GL-$g$ ($\tau^{2}=1$)”), one using a hyper-$g/n$ prior (labeled “GL-$g$ ($\tau^{2}=n$)”), and one using a scaled hyper-$g$ prior where the scale is given a half-Cauchy hyperprior (labeled “GL-$g$ ($\tau^{2}\sim\mathsf{HC}$) ”); (e) the adaptive Lasso (ALasso) of (Huang et al., 2008); (f) the Horseshoe prior (Carvalho et al., 2010, labeled “Horseshoe” in the sequel) and (g) the Horseshoe-Pit prior (Denti et al., 2023, labeled “HSM”). Computation under standard mixtures of $g$ priors relies on version 1.7.1 of the R package BAS, while computation under block $g$ priors with known blocking structures relies on a slight variation of our own code for the Diriclet mixtures of $g$ priors. Computation for the adaptive Lasso relies on version 4.1.6 of the R package glmnet. Computation under the Horseshoe prior relies on version 1.2 of the R package bayesreg, while computation under the Horseshoe-Pit prior relies on code from the author of that manuscript which, at the time of this writing, is available at https://github.com/Fradenti/HorseshoeMix. For all Bayesian procedures that require a prior on model space, we assign $\boldsymbol{\gamma}$ a Beta-Binomial prior with $c=d=1$ (e.g., see Scott and Berger, 2010 and Porwal and Raftery, 2022 and Section 5 above). Furthermore, in order to avoid improper priors when $p\geq n$, the prior on models is constrained so that models for which $\sum_{k=1}^{p}\gamma_{i}>n-2$ receive zero probability.
Power (large coeffs) Power (small coeffs) Type I error (null coeffs) Power (large coeffs) Power (small coeffs) Type I error (null coeffs)
$p=250$
$\eta=0$ $\eta=0.9$
g-prior ($\tau^{2}=$n) 0.986 0.856 0.005 0.949 0.506 0.004
Som et al. ($\tau^{2}=$n, K=2) 0.986 0.913 0.037 0.962 0.850 0.219
Som et al. ($\tau^{2}=$n, K=3) 0.992 1.000 1.000 0.980 1.000 1.000
GL-$g$ ($\tau^{2}=$1) 1.000 1.000 1.000 1.000 1.000 1.000
GL-$g$ ($\tau^{2}=$n) 0.990 0.908 0.054 0.969 0.694 0.046
GL-$g$ ($\tau^{2}\simHC$) 0.990 0.900 0.039 0.973 0.745 0.100
DP block-g ($\tau^{2}=$1) 0.991 0.921 0.160 0.984 0.843 0.394
DP block-g ($\tau^{2}=$n) 0.989 0.904 0.045 0.974 0.760 0.126
DP block-g ($\tau^{2}\simHC$) 0.990 0.906 0.049 0.977 0.788 0.211
ALasso 0.956 0.569 0.011 0.873 0.352 0.157
Horseshoe 0.989 0.889 0.027 0.965 0.656 0.022
HSM 0.986 0.860 0.005 0.962 0.625 0.011
$p=500$
$\eta=0$ $\eta=0.9$
g-prior ($\tau^{2}=$n) 0.979 0.814 0.000 0.913 0.342 0.001
Som et al. ($\tau^{2}=$n, K=2) 0.982 0.873 0.005 0.944 0.636 0.028
Som et al. ($\tau^{2}=$n, K=3) 0.986 0.923 1.000 0.963 0.843 1.000
GL-$g$ ($\tau^{2}=$1) 1.000 1.000 0.993 0.999 0.992 0.973
GL-$g$ ($\tau^{2}=$n) 0.984 0.866 0.006 0.955 0.580 0.010
GL-$g$ ($\tau^{2}\simHC$) 0.984 0.865 0.005 0.957 0.597 0.014
DP block-g ($\tau^{2}=$1) 0.985 0.873 0.014 0.961 0.636 0.040
DP block-g ($\tau^{2}=$n) 0.984 0.867 0.006 0.957 0.612 0.016
DP block-g ($\tau^{2}\simHC$) 0.984 0.867 0.007 0.958 0.622 0.023
ALasso 0.745 0.117 0.056 0.269 0.008 0.006
Horseshoe 0.982 0.842 0.012 0.947 0.540 0.007
HSM 0.981 0.831 0.001 0.943 0.506 0.001
$p=750$
$\eta=0$ $\eta=0.9$
g-prior ($\tau^{2}=$n) 0.959 0.607 0.000 0.887 0.212 0.000
Som et al. ($\tau^{2}=$n, K=2) 0.982 0.863 0.003 0.941 0.576 0.015
Som et al. ($\tau^{2}=$n, K=3) 0.985 0.884 0.488 0.957 0.703 0.563
GL-$g$ ($\tau^{2}=$1) 0.992 0.922 0.468 0.976 0.792 0.464
GL-$g$ ($\tau^{2}=$n) 0.985 0.849 0.003 0.946 0.506 0.005
GL-$g$ ($\tau^{2}\simHC$) 0.985 0.851 0.003 0.948 0.526 0.008
DP block-g ($\tau^{2}=$1) 0.985 0.857 0.005 0.952 0.547 0.014
DP block-g ($\tau^{2}=$n) 0.985 0.852 0.003 0.949 0.536 0.010
DP block-g ($\tau^{2}\simHC$) 0.985 0.853 0.004 0.951 0.548 0.013
ALasso 0.722 0.106 0.054 0.242 0.007 0.004
Horseshoe 0.980 0.800 0.005 0.940 0.456 0.004
HSM 0.982 0.818 0.000 0.938 0.449 0.001
Table 1: Estimates of power for small (generated from a $\mathcal{N}(0,1)$ distribution) and large (generated from a $\mathcal{N}(0,10)$ distribution) coefficients, and of type I error for null coefficients ($\beta=0$) in our second simulation study. For the purpose of this table, coefficients are considered “significant” is their posterior inclusion probability is greater than 0.5.
In the introduction, we motivated DP mixtures of block $g$ priors by arguing that it should lead to higher power for detecting smaller coefficients. We also claimed that procedures that pre-select a blocking of the coefficients can be very sensitive to this choice. Evidence of these claims is presented in Table 1, which shows the estimated power associated with identifying “large” (those generated from a zero-mean normal with standard deviation 10) and “small” (those generated from a standard normal distribution) coefficients, as well as the type I error (for the null coefficients) under the various procedures. For the purpose of this table, coefficients are considered “significant” if their posterior inclusion probability is greater than 0.5 (in the case of procedures based on variants of the $g$-prior), if they are included in the optimal model after applying generalized cross-validation to identify the optimal penalty parameter (for ALasso), or if their 95% posterior credible intervals do not cover zero (in the case of Horseshoe and HSM). First, we note that ALasso shows by far the lowest power to detect small coefficients. Even for large coefficients, the performance of ALasso degrades substantially in “large pp” scenarios, especially when variables are highly correlated. The same is true of other penalized likelihood procedures we tried (results not shown). Next, we note that the standard $g$ prior tends to have lower power than the other Bayesian procedures we consider when it comes to detecting small coefficients. Indeed, somewhat surprisingly, HSM and Horseshoe seem to have higher overall power than standard $g$-priors in this scenario with similar type-I errors, even though they were not originally designed for model selection. GL-$g$ procedures with $\tau^{2}=1$ also seem to perform poorly, yielding high power for detecting small signals, but also very high type I error rates. Similarly to GL-$g$ with $\tau^{2}=1$, the procedure of Som (2014) with $K=3$ tends to have the highest power, but it also comes with extremely high type I error rates. This is because, by parceling out the null coefficients into a separate cluster, the block $g$ prior ends up overfitting by learning a very small shrinkage factor for the coefficients in this block. The results for Som (2014) with $K=2$ indicate that this issue can be addressed by having the null coefficients assigned to blocks that contain some significant coefficients, avoiding overfitting. This solution is, however, clearly impractical in real applications, as we do not know which coefficients are likely to be significant in the first place. The best performing procedures correspond to GL-$g$ ($\tau^{2}=n$), GL-$g$ ($\tau^{2}\sim\mathsf{HC}$), and the three procedures based on Dirichlet process mixtures of block $g$-priors. These are all priors that can be understood as being either “unit information”, or allowing for enough flexibility to learn the expected value of the $g_{j}$s from the data. For $\eta=0$, all of these 5 procedures seem to yield very similar performance in terms of power and type I errors. However, for $\eta=0.9$, procedures based on DP block-$g$ priors consistently show higher power for detecting small coefficients, at the price of a very small increase (in absolute terms) in the type I error.
Figure 3: $F_{1}$ scores for model selection procedures for our second simulation study. These $F_{1}$ scores excludes the large coefficients in the calculation of precision and recall.
To investigate the false positive/false negative tradeoffs associated with the various methods, we present in Figure 3 boxplots across the various simulated datasets of $F_{1}$ for the various procedures. Recall that the $F_{1}$ score is defined as the harmonic mean of proportion of true positives among “selected” covariates (the precision) and the proportion of “selected” covariates among true positive covariates (the recall). The $F_{1}$ score ranges between 0 and 1, with a higher value indicating better model selection performance. As would be expected from our discussion of power and false positive rates, ALasso is by far the worst performer, followed by Som et al. $(K=3)$ and procedures based on the GL-$g$ ($\tau^{2}=1$) prior. Procedures based on the standard $g$-prior perform somewhat better than these three, but they are still suboptimal, particularly when the covariates are highly correlated. The performance of the remaining procedures is very similar, but some subtle patterns are still visible. Som et al. $(K=2)$ seems to perform slightly better than the rest but, as discussed before, it relies on the unrealistic assumption that we are able to distinguish between large and small coefficients up front. Similarly, GL-$g$ ($\tau^{2}=n$), Horseshoe and HSM seem to slightly underperform, particularly when the correlation is high.
(a) $\eta=0$
(b) $\eta=0.9$
Figure 4: Normalized prediction MSE for $\eta=0$ and $\eta=0.9$. Normalization is with respect to the prediction MSE under the standard $g$-prior.
Taken together, the results in Table 1 and Figure 3 highlight (a) the benefits of using differential shrinkage priors in the context of model selection, (b) the need to either properly center the distribution of the shrinkage coefficients or, alternatively, allowing enough model flexibility to learn its center from its data, and (c) the risks associated with the use of fixed rather than data-driven blocks in the development of model selection priors.
Finally, we present in Figure 4 the prediction mean square error (MSE) for each of the procedures. To obtain these prediction MSEs, each dataset was augmented with a test set of 500 additional observations. Furthermore, in order to simply interpretation, we normalized all MSEs with respect to that under the $g$-prior for each dataset and graph the resulting ratios in a logarithmic scale. This means that values less than 1 correspond to methods with smaller (better) prediction MSE. Note that, with the exception of ALasso, all procedures consistently outperform $g$-priors in “large pp” regimes, again illustrating the the advantages of differential shrinkage in Bayesian settings. Note that Horseshoe has worse performance than the other Bayesian procedures, particularly in “large pp” scenarios. While perhaps surprising at first sight, this observation is consistent with the results of Lee et al. (2020). Som et al. $(K=3)$ seems to perform slightly better than Som et al. $(K=2)$ in this evaluation, specially when $\eta=0.9$, which is the opposite of what we observed when evaluating model selection performance. Similar to model selection performance, procedures based of the GL-$g$ ($\tau^{2}=1$) prior also have worse performance than the remaining GL-$g$ and DP block-$g$ procedures, all of which perform quite similarly in terms of predictions. Additional results related to simulation study, including results for $\eta=0.5$, examples of posterior distributions over $(p_{\boldsymbol{\gamma}},K_{\boldsymbol{\gamma}})$, and a sensitivity analysis for priors on the concentration parameter $\alpha$ and model space $\boldsymbol{\gamma}$ can be seen in the supplementary materials.
7 The ozone dataset
We further investigate the performance of DP mixtures of block $g$ prior using the ozone dataset introduced in Breiman and Friedman (1985) and later analyzed in Casella and Moreno (2006) and Liang et al. (2008), among others. The dataset consists of daily measurements of the maximum ozone concentration near Los Angeles and eight meteorological variables. We consider regression models that might include all eight of these variables along with all possible interactions and squares, leading to up to 44 possible predictors.
(a) Posterior inclusion probabilities
(b) Model size
Figure 5: Posterior inclusion probabilities for individual variables and model sizes for various model selection procedures in the ozone dataset.
Figure 5(a) shows the posterior inclusion probabilities (PIPs) for each of the predictors (i.e., $\mathsf{Pr}(\gamma_{i}=1\mid\boldsymbol{y})$) under consideration for the various competing procedures described in Section 6.2. In the case of ALasso, these are taken to be 1 if the variable is non-zero in the model fitted using optimal penalty parameter according to generalized crossvalidation. On the other hand, for Horseshoe and HSM, the PIPs are reported as 0 if the 95% posterior credible interval for the variable includes 0, and as 1 otherwise. Note that ALasso is an outlier and tends to select a much larger number of variables (17) than any of the Bayesian procedures. On the other hand, there is fair bit of agreement in the PIPs among the various Bayesian procedures. For example, all of them agree in that sbtp (Sandburg Air Force Base temperature) and ibht (inversion base height at LAX) should be included in the model. There are, however, interesting differences as well. For example, HSM, Horseshoe and the standard $g$-prior all agree in including the square of hmdt (humidity) in the model, but not the main effect of hmdt. In contrast, GL-$g$ and DP block-$g$ assign moderate probabilities of inclusion to both the linear and quadratic terms associated with humidity instead. To complement these results, we show in Figure 5(a) the posterior distribution of $p_{\boldsymbol{\gamma}}$, the number of variables included in the model, for DP block-$g$, GL-$g$ and the standard $g$-prior. Interestingly, the standard $g$ prior tends to include the most variables (in some cases, as many as ALasso), while GL-$g$ tends to selects the most parsimonious models. As would be expected, DP block-$g$ is somewhere in between them.
Figure 6: Joint and marginal posterior distributions for $p_{\gamma}$ and $K_{\boldsymbol{\gamma}}$ under the DP mixtures of block $g$ prior for the ozone dataset.
To gain additional insight into the behavior of the various approaches, Figure 6 shows the joint and marginal posterior distributions for $p_{\gamma}$ and $K_{\boldsymbol{\gamma}}$ (the number of blocks in which the $p_{\gamma}$ included variables have been grouped) under the DP mixture of block $g$ priors. Note that the number of variables included by this procedure ranges between 4 and 17, with a clear mode at 7. The procedure also places moderate probability (around 0.49) on models that group these variable into more than one block of variables, but virtually no probability to any model with more than 8 or 9 blocks. This result is consistent with our previous observation that procedures based on DP block-$g$ adaptively “interpolate” between those produced by standard $g$ priors and those generate by GL-$g$.
(a) Predictive mean squared error
(b) Median Interval Scores
Figure 7: Predictive mean squared error (MSE) and and median interval scores (MIS) for our crossvalidation exercise for the ozone dataset. Note that MIS is not readily available for ALasso or other penalized likelihood methods, so it is not included.
Finally, Figure 7 presents boxplots of the predictive mean squared error (MSE) and 95% median intervals scores (MIS, Gneiting and Raftery, 2007) for a crossvalidation exercise in which 20 random 80-20 splits of the data were used to train and then test prediction accuracy. For a variable $z$, the $\alpha\times 100\%$ IS is given by $IS_{\alpha}(l,u,z)=(u-l)+\frac{2}{1-\alpha}(l-z)\mathbbm{1}\{z<l\}+\frac{2}{1-\alpha}(z-u)\mathbbm{1}\{u<z\}$, where $l$ and $u$ denote the upper and lower bounds of the $\alpha\times 100\%$ posterior intervals of $z$. The first term in this expression rewards narrow predictive intervals, while the second rewards accurate coverage. We do not report the MIS for ALasso because the implementation in the glmnet package does not provide a measure of predictive uncertainty off the shelf. Generally speaking, Horseshoe, HSM and ALasso seem to have a slightly better predictive performance than procedures based on standard $g$-priors, GL-$g$ priors and DP block $g$ priors, particularly for point prediction. However, the differences are small.
8 Discussion
We introduced novel class of priors that that enable a parsimonious, data driven approach to model selection and prediction in linear models that is free from the so-called conditional Lindley “paradox” and that provides a bridge between two strands of the literature (model selection priors and continuous shrinkage priors) that have often been treated as distinct. The use of a Dirichlet process prior for the distribution of the local shrinkage coefficients enable us to interpret our model in terms of blocks of covariates that receive the same level of shrinkage. However, the method is not specially reliant to the clustering properties of Dirichlet process. In particular, it does not require that we are able to consistently estimate the “true” partition of the coefficients, only that the partitions that receive high probability a posteriori be those that do not mix both “small” and “large” coefficients.
While this paper focuses on linear regression models, DP mixtures of block $g$ priors can be extended to generalized linear models, and perhaps even non-linear regression, by setting $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}$ to be an appropriate information matrix, e.g., see Bové and Held (2011), Li and Clyde (2018) and Porwal and Rodríguez (2023). The approach introduced here can also be used to generalize the class of priors introduced in Carvalho and Scott (2009), leading to a new class of (mixtures of) hyper-inverse Wishart block $g$ priors for model selection in Gaussian graphical models.
Recently, Lee et al. (2024) demonstrated that, in the context of continuous shrinkage priors, the behavior of the tail of the distribution of the local shrinkage parameters plays an important role in the performance of the prior, and that different types of tail behavior might be necessary in sparse vs. ultra-sparse settings. The use of a non-parametric specification for the distribution of $g_{j}\mid\boldsymbol{\gamma}$ through a Dirichlet process prior enables such tail adaptability in the context of model selection priors.
From a theoretical perspective, there are two additional aspects of our work that are open for extension. First, our results around the conditional Lindley paradox assume that the design matrix is orthogonal. The evidence from our simulations suggested that the results, and in particular the ability of the model to separate “large” and “small” coefficients into separate clusters, extends to the non-orthogonal case. We believe that theoretical progress in this area can be achieved by developing an asymptotic expansion for the multivariate integral defining the marginal likelihood. Similarly, our model consistency results assume that $p$ is fixed. We believe that it possible to extend the result to the cases where $p$ grows with $n$. Both of these directions will be explored elsewhere.
9 Supplementary Materials
Supplementary materials contain the derivation of marginal likelihood in Equation (3.2), visualization of the tail behavior of the prior for the bivariate case, proof of theorems in Section 4, details of the MCMC algorithm discussed in Section 5 and additional simulation results for Section 6. An implementation of the MCMC algorithm is available at https://github.com/Anupreet-Porwal/DP-mix-block-g-prior.
10 Acknowledgment
We are grateful to the editor, the associate editor, and the referees for their valuable and constructive comments on an earlier version of this article. We also thank Prof. Merlise Clyde for her insights during the early stages of the project.
11 Disclosure Statement
The authors report there are no competing interests to declare.
12 Funding
This work was partially supported by grants NSF-2023495, NSF-2114727 and NSF-2523615.
Appendix A Derivation of Equation (3.2)
By definition,
$f(\boldsymbol{y}\mid\boldsymbol{\gamma},g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}})=\int f(\boldsymbol{y}\mid\boldsymbol{\beta}_{\boldsymbol{\gamma}},\sigma^{2})f(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\sigma^{2},\boldsymbol{G}_{\boldsymbol{\gamma}},\boldsymbol{\gamma})f(\beta_{0},\sigma^{2})\mathop{}\!\mathrm{d}\beta_{0}\mathop{}\!\mathrm{d}\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mathop{}\!\mathrm{d}\sigma^{2}$
The integral with respect to $\boldsymbol{\beta}_{\boldsymbol{\gamma}}$ is trivial to compute using the properties of the multivariate normal distribution, resulting in $\boldsymbol{y}\mid\beta_{0},\sigma^{2},\boldsymbol{\gamma},g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}}\sim\mathsf{N}\left(\boldsymbol{1}\beta_{0},\sigma^{2}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}\right)$, where $\boldsymbol{\Omega}_{\boldsymbol{\gamma}}=\boldsymbol{I}_{n}+\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}$. Integrating now with respect to $\beta_{0}$
$f(\boldsymbol{y}\mid\boldsymbol{\gamma},\sigma^{2},g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}})$ $=\int(2\pi\sigma^{2})^{-n/2}\left|\boldsymbol{\Omega}_{\boldsymbol{\gamma}}\right|^{-1/2}$
$\;\;\;\;\;\;\;\;\;\exp\left\{-\frac{1}{2\sigma^{2}}\left(\boldsymbol{y}-\boldsymbol{1}\beta_{0}\right)^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\left(\boldsymbol{y}-\boldsymbol{1}\beta_{0}\right)\right\}\mathop{}\!\mathrm{d}\beta_{0}$
$=(2\pi\sigma^{2})^{-n/2}\left|\boldsymbol{\Omega}_{\boldsymbol{\gamma}}\right|^{-1/2}\exp\left\{-\frac{1}{2\sigma^{2}}\left(\boldsymbol{y}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{y}-\frac{\left(\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{y}\right)^{2}}{\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{1}_{n}}\right)\right\}$
$\;\;\;\;\;\;\;\;\;\int\exp\left\{-\frac{\boldsymbol{1}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{1}}{2\sigma^{2}}\left(\beta_{0}-\frac{\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{y}}{\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{1}_{n}}\right)^{2}\right\}\mathop{}\!\mathrm{d}\beta_{0}$
$=\left(\frac{1}{2\pi\sigma^{2}}\right)^{\frac{n-1}{2}}\frac{\left|\boldsymbol{\Omega}_{\boldsymbol{\gamma}}\right|^{-1/2}}{\left(\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\gamma}^{-1}\boldsymbol{1}_{n}\right)^{1/2}}$
$\;\;\;\;\;\;\;\;\exp\left\{-\frac{1}{2\sigma^{2}}\left(\boldsymbol{y}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{y}-\frac{\left(\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{y}\right)^{2}}{\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{1}_{n}}\right)\right\}.$
Finally,
$f(\boldsymbol{y}\mid\boldsymbol{\gamma},\sigma^{2},g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}})$ $=\left(\frac{1}{2\pi}\right)^{\frac{n-1}{2}}\frac{\left|\boldsymbol{\Omega}_{\boldsymbol{\gamma}}\right|^{-1/2}}{\left(\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\gamma}^{-1}\boldsymbol{1}_{n}\right)^{1/2}}$
$\;\;\;\;\;\;\;\;\int\left(\frac{1}{\sigma^{2}}\right)^{\frac{n-1}{2}}\exp\left\{-\frac{1}{2\sigma^{2}}\left(\boldsymbol{y}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{y}-\frac{\left(\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{y}\right)^{2}}{\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{1}_{n}}\right)\right\}\frac{1}{\sigma^{2}}\mathop{}\!\mathrm{d}\sigma^{2}$
$=\frac{\Gamma\left(\frac{n-1}{2}\right)}{\pi^{\frac{n-1}{2}}}\frac{\left|\boldsymbol{\Omega}_{\boldsymbol{\gamma}}\right|^{-1/2}}{\left(\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\gamma}^{-1}\boldsymbol{1}_{n}\right)^{1/2}}\left[\boldsymbol{y}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{y}-\frac{\left(\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{y}\right)^{2}}{\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{1}_{n}}\right]^{-\frac{n-1}{2}}.$
Note, however, that since the design matrix has been centered, $\boldsymbol{1}_{n}^{T}\boldsymbol{X}=\boldsymbol{0}$. This implies $\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\gamma}^{-1}\boldsymbol{1}_{n}=n$ and $\boldsymbol{1}_{n}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{y}=\sum_{i=1}^{n}y_{i}$, and yields the simplified form in (3.2).
Appendix B Scatterplots of realizations of various prior distributions in the bivariate case
Figure 8 shows scatterplots of random samples from the Dirchlet mixture of block $g$ priors and some related distributions in the bivariate case under a hyper-$g/n$ distribution. In this special case, the DP block $g$ prior in panel (d) is a mixture of the standard $g$ prior in panel (a) and the block $g$ prior in panel (c). Note that the contour plots for (8b), (8c) and (8d) are non-elliptical, unlike those of (8a). These non-elliptical contours indicate that these priors allow for direction-dependent shrinkage, which is what allows them to avoid the conditional Lindley paradox. Furthermore, note that the axes of symmetry of the contour plots of the global-local $g$ prior and the DP block $g$ prior are not parallel to any of the main axes, a consequence of the fact that both priors account for prior correlation among the explanatory variables (unlike the orthogonal block $g$ prior of Som et al., 2016).
(a) Standard $g$ prior
(b) Orthogonal block $g$ prior
(c) Global-local $g$ prior
(d) DP mixture of block $g$ priors
Figure 8: Scatterplots of random samples from the Dirchlet mixture of block $g$ priors and some related distributions in the bivariate case under a hyper-$g/n$ distributon for the shrinkage parameter(s). Panel (a) corresponds to the (elliptical) contours of the standard $g$ prior of Liang et al. (2008). Panel (b) shows the density of the prior proposed by Som et al. (2016), which assumes that blocks are orthogonal a priori. Panel (c) corresponds to a global-local $g$ prior in which each covariate is assigned is own independent shrinkage factor and the prior covariance matrix is proportional to $\left(\boldsymbol{X}^{T}\boldsymbol{X}\right)^{-1}$. Panel (d) is our DP mixture of block $g$ priors, which in this case corresponds to a mixture of the distributions in panels (a) and (c).
Appendix C Proof of Theorems 4.1 and 4.2
Two results will be useful in what follows. First, the Woodbury matrix identity implies
$\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}=\left[\mathbf{I}_{n}+\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\left(\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right)^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\right]^{-1}=\\ \mathbf{I}_{n}-\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\left(\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right)^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}.$
Secondly, using the matrix determinant lemma
$\left|\boldsymbol{\Omega}_{\boldsymbol{\gamma}}\right|=\left|\mathbf{I}_{n}+\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\left(\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right)^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\right|=\frac{\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right|}{\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right|}.$
Now, on to the proof of Theorem 4.1. From equation (3.2), we have
$\frac{f(\boldsymbol{y}\mid\boldsymbol{\gamma},\boldsymbol{G})}{f(\boldsymbol{y}\mid\boldsymbol{\gamma}=\boldsymbol{0})}$ $=\left|\boldsymbol{\Omega}_{\boldsymbol{\gamma}}\right|^{-1/2}\left[\frac{\boldsymbol{y}^{T}\boldsymbol{y}-n\bar{y}^{2}}{\boldsymbol{y}^{T}\boldsymbol{\Omega}_{\boldsymbol{\gamma}}^{-1}\boldsymbol{y}-n\bar{y}^{2}}\right]^{\frac{n-1}{2}}$
$=\frac{\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right|^{1/2}}{\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right|^{1/2}}$
$\;\;\;\;\left[\frac{\boldsymbol{y}^{T}\boldsymbol{y}-n\bar{y}^{2}}{\boldsymbol{y}^{T}\boldsymbol{y}-n\bar{y}^{2}-\boldsymbol{y}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\left(\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right)^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{y}}\right]^{\frac{n-1}{2}}$
$=\frac{\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right|^{1/2}}{\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right|^{1/2}}$
$\;\;\;\;\left[1-R^{2}_{\boldsymbol{\gamma}}\frac{\boldsymbol{y}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\left(\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right)^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{y}}{\boldsymbol{y}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\left(\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right)^{-1}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{y}}\right]^{-\frac{n-1}{2}}$
$=\frac{\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right|^{1/2}}{\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right|^{1/2}}$
$\;\;\;\;\left[1-R^{2}_{\boldsymbol{\gamma}}\frac{\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}^{T}\left(\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}-\left[\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\}^{-1}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right]^{-1}\right)\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}}{\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}}\right]^{-\frac{n-1}{2}}$
where $\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}$ is the maximum likelihood estimator under model $\boldsymbol{\gamma}$ and $R^{2}_{\boldsymbol{\gamma}}$ is the coefficient of determination. Now, let
$\Upsilon(\boldsymbol{G}_{\boldsymbol{\gamma}},\boldsymbol{y})=\frac{\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}^{T}\left(\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}-\left[\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\}^{-1}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right]^{-1}\right)\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}}{\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}}.$
Then, the condition required for the Bayes factor to be information consistent in this setting can be written as
$\lim_{\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}\|\to\infty}\int\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right|^{-1/2}\\ \left[1-R^{2}_{\boldsymbol{\gamma}}\Upsilon(\boldsymbol{G}_{\boldsymbol{\gamma}},\boldsymbol{y})\right]^{-\frac{n-1}{2}}f(g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}})\mathop{}\!\mathrm{d}g_{1}\cdots\mathop{}\!\mathrm{d}g_{p_{\boldsymbol{\gamma}}}=\infty$
which, because of dominated convergence, can be written as
$\int\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right|^{-1/2}\\ \left\{\lim_{\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}\|\to\infty}\left[1-R^{2}_{\boldsymbol{\gamma}}\Upsilon(\boldsymbol{G}_{\boldsymbol{\gamma}},\boldsymbol{y})\right]^{-\frac{n-1}{2}}\right\}f(g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}})\mathop{}\!\mathrm{d}g_{1}\cdots\mathop{}\!\mathrm{d}g_{p_{\boldsymbol{\gamma}}}=\infty$ (C.1)
Recalling that $\lim_{\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}\|\to\infty}R^{2}_{\boldsymbol{\gamma}}=1$, (C.1) reduces to
$\int\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right|^{-1/2}\\ \left[1-\Upsilon^{*}(\boldsymbol{G}_{\boldsymbol{\gamma}})\right]^{-\frac{n-1}{2}}f(g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}})\mathop{}\!\mathrm{d}g_{1}\ldots\mathop{}\!\mathrm{d}g_{p_{\boldsymbol{\gamma}}}=\infty.$
where $\Upsilon^{*}(\boldsymbol{G}_{\boldsymbol{\gamma}})=\lim_{\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}\|\to\infty}\Upsilon(\boldsymbol{G}_{\boldsymbol{\gamma}},\boldsymbol{y})$.
Now, since $\boldsymbol{X}$ is full rank and all $g_{j}$s are strictly positive, both $\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}$ and $\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}-\left[\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\}^{-1}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right]^{-1}$ are strictly positive definite matrices. It follows then that
$0<\lambda_{-}(\boldsymbol{G})\left\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}\right\|^{2}\leq\\ \hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}^{T}\left(\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}-\left[\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\}^{-1}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right]^{-1}\right)\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}\leq\\ \lambda_{+}(\boldsymbol{G})\left\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}\right\|^{2}<\infty,$
and
$0<\nu_{-}\left\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}\right\|^{2}\leq\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}^{T}\boldsymbol{X}\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}\leq\nu_{+}\left\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}\right\|^{2}<\infty,$
where $\lambda_{-}(\boldsymbol{G})$ and $\lambda_{+}(\boldsymbol{G})$ are the largest and the smallest eigenvalues of the matrix $\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}-\left[\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\}^{-1}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right]^{-1}$ and $\nu_{+}$ and $\nu_{-}$ are the largest and smallest eigenvalues of $\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}$. Therefore
$0<\frac{\lambda_{-}(\boldsymbol{G})}{\nu_{+}}\leq\Upsilon(\boldsymbol{G}_{\boldsymbol{\gamma}},\boldsymbol{y})\leq\frac{\lambda_{+}}{\nu_{-}}(\boldsymbol{G})<\infty.$
These bounds are independent of $\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}$ and therefore apply to $\Upsilon^{*}(\boldsymbol{G}_{\boldsymbol{\gamma}})$ as well. Hence:
$\int\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right|^{-1/2}\left[1-\Upsilon^{*}(\boldsymbol{G}_{\boldsymbol{\gamma}})\right]^{-\frac{n-1}{2}}f(g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}})\mathop{}\!\mathrm{d}g_{1}\ldots\mathop{}\!\mathrm{d}g_{p_{\boldsymbol{\gamma}}}\geq\\ \int\left|\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}+\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right|^{-1/2}\left[1-\frac{\lambda_{-}(\boldsymbol{G})}{\nu_{+}}\right]^{-\frac{n-1}{2}}f(g_{1},\ldots,g_{p_{\boldsymbol{\gamma}}})\mathop{}\!\mathrm{d}g_{1}\ldots\mathop{}\!\mathrm{d}g_{p_{\boldsymbol{\gamma}}}.$
This completes the proof of Theorem 4.1.
The proof of Theorem 4.2 relies on the fact that $BF_{\boldsymbol{\gamma},\boldsymbol{0}}(\boldsymbol{y})$ under the Dirichlet mixtures of block $g$ prior can be written as a weighted average of Bayes factors conditional on each of the possible partitions of the $p_{\boldsymbol{\gamma}}$ coefficients,
$BF_{\boldsymbol{\gamma},\boldsymbol{0}}(\boldsymbol{y})=\sum_{\rho}p(\rho)BF_{\boldsymbol{\gamma},\boldsymbol{0}}(\boldsymbol{y}\mid\rho)$ (C.2)
where $p(\rho)=\int p(\rho\mid\alpha)p(\alpha)\mathop{}\!\mathrm{d}\alpha$. Note that one of the terms in the sum corresponds to $\rho=\rho_{0}=\left\{\{1,2,3,\ldots,p\}\right\}$, i.e., the Bayes factor under the standard $g$ prior. Hence, if the density of the base measure satisfies
$\int(1+g)^{(n-1-p_{\boldsymbol{\gamma}})/2}f(g\mid\tau^{2},a,b)\mathop{}\!\mathrm{d}g=\infty$
then we know that
$\lim_{\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}\|\to\infty}BF_{\boldsymbol{\gamma},\boldsymbol{0}}(\boldsymbol{y}\mid\rho_{0})=\infty.$
But all the other conditional Bayes factors $BF_{\boldsymbol{\gamma},\boldsymbol{0}}(\boldsymbol{y}\mid\rho)$ in (C.2) are non-negative, so we must have $\lim_{\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}\|\to\infty}BF_{\boldsymbol{\gamma},\boldsymbol{0}}(\boldsymbol{y})=\infty$.
Appendix D Proof of Theorem 4.3
We start by introducing some notation. Let $\rho=\{S_{1},\ldots,S_{K}\}$ be a partition of $\mathcal{I}$, $m_{k}=|S_{k}|$ be the number of elements in $S_{k}$ and, similarly, $m_{1,k}=|S_{k}\cap\mathcal{I}_{1}|$ and $m_{2,k}=|S_{k}\cap\mathcal{I}_{2}|$. Clearly, $m_{1,k},m_{2,k}\geq 0$ and $m_{1,k}+m_{2,k}=m_{k}$.
Consider first the case where $\sigma^{2}$ is known. Because $\boldsymbol{X}$ is orthogonal, it is easy to verify that, under (3.3),
$f\left(\boldsymbol{y}(N)\mid\rho,\sigma^{2}\right)$ $\propto\left(\frac{1}{\sigma^{2}}\right)^{n/2}\prod_{k=1}^{K}\int_{0}^{1}u_{k}^{b}(1-u_{k})^{a+\frac{m_{k}}{2}}\exp\left\{\frac{u_{k}\|\hat{\boldsymbol{\beta}}_{S_{k}}(N)\|}{2\sigma^{2}}\right\}\mathop{}\!\mathrm{d}u_{k}$
$\propto\left(\frac{1}{\sigma^{2}}\right)^{n/2}\prod_{k=1}^{K}M\left(b+1,a+b+\frac{m_{k}}{2}+2,\frac{\|\hat{\boldsymbol{\beta}}_{S_{k}}(N)\|}{2\sigma^{2}}\right)$
where $M$ is Kummer’s function (Abramowitz et al., 1988), $\|\hat{\boldsymbol{\beta}}_{S_{k}}(N)\|=\sum_{j\in S_{k}}\hat{\beta}_{j}^{2}(N)$, and $\hat{\beta}_{j}(N)$ is the maximum likelihood estimator of $\beta_{j}$ based on $\boldsymbol{y}(N)$. A well known asymptotic expansion of Kummer’s function is $M(a,b,z)\approx\frac{\Gamma(b_{0})}{\Gamma(a_{0})}z^{a_{0}-b_{0}}\exp\{z\}$ for large $z$ (see Equation 13.5.1 in page 508 of Abramowitz et al., 1988). Hence, under the assumptions of the theorem,
$f\left(\boldsymbol{y}(N)\mid\rho,\sigma^{2}\right)$ $\approx\exp\left\{\mathcal{O}(N)\right\}\mathcal{O}\left(N^{1-a-\sum_{\{k:m_{1,k}>0\}}m_{k}/2}\right)$
for large $N$. Now, if $\rho\preceq\rho_{0}$, then $\sum_{\{k:m_{1,k}>0\}}|S_{k}|=p_{1}$ and therefore
$\lim_{N\to\infty}\frac{f\left(\boldsymbol{y}(N)\mid\rho,\sigma^{2}\right)}{f\left(\boldsymbol{y}(N)\mid\rho_{0},\sigma^{2}\right)}=\lim_{N\to\infty}\frac{\exp\left\{\mathcal{O}(N)\right\}\mathcal{O}\left(N^{1-a-p_{1}/2}\right)}{\exp\left\{\mathcal{O}(N)\right\}\mathcal{O}\left(N^{1-a-p_{1}/2}\right)}=c_{\rho}$
for some $0<c_{\rho}<\infty$.
On the other hand, if $\rho\npreceq\rho_{0}$ then there exists at least one $S_{k}$ such that both $m_{1,k}>0$ and $m_{2,k}>0$. Recall that $m_{1,k}+m_{2,k}=m_{k}$. Therefore, for a partition $\rho\npreceq\rho_{0}$, we have $p_{1}=\sum_{k}m_{1,k}=\sum_{\{k:m_{1,k}>0\}}m_{1,k}<\sum_{\{k:m_{1,k}>0\}}m_{k}$. Hence, in this case,
$\lim_{N\to\infty}\frac{f\left(\boldsymbol{y}(N)\mid\rho,\sigma^{2}\right)}{f\left(\boldsymbol{y}(N)\mid\rho_{0},\sigma^{2}\right)}=0.$
When $\sigma^{2}$ is unknown, note that Som (2014) shows that the limit as $N\to\infty$ of the posterior distribution for $\sigma^{2}$ conditional on the partition is a proper, non-degenerate distribution. Hence, we have the asymptotic expansion in this case is instead
$f\left(\boldsymbol{y}(N)\mid\rho\right)$ $\approx\mathcal{O}\left(N^{1-a-\sum_{\{k:m_{1,k}>0\}}m_{k}/2}\right).$
for large $N$. Hence, we again have
$\lim_{N\to\infty}\frac{f\left(\boldsymbol{y}(N)\mid\rho\right)}{f\left(\boldsymbol{y}(N)\mid\rho_{0}\right)}=\begin{cases}0&\rho\nprec\rho_{0},\\ c_{\rho}&\rho\prec\rho_{0}\end{cases}$
for come $0<c_{\rho}<\infty$.
Appendix E Proof of Theorem 4.4
As in Theorem 4.3, let $\mathcal{I}_{1}=\{j^{(1)}_{1},\ldots,j^{(1)}_{p_{1}}\}$ denote the set of indexes associated with the covariates included in the design matrix $\boldsymbol{X}_{1}$, $\mathcal{I}_{2}=\{j^{(2)}_{1},\ldots,j^{(2)}_{p_{2}}\}$ be the set associated with $\boldsymbol{X}_{2}$, and $\mathcal{I}=\mathcal{I}_{1}\cup\mathcal{I}_{2}$. Also, recall that $\mathcal{I}_{1}\cap\mathcal{I}_{2}=\emptyset$. Now, note that
$BF_{\boldsymbol{\gamma}_{a},\boldsymbol{\gamma}_{0}}(\boldsymbol{y}(N))$ $=\frac{\sum_{\rho}f(\rho\mid\boldsymbol{\gamma}_{a})f\left(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{a},\rho\right)}{\sum_{\rho}f(\rho\mid\boldsymbol{\gamma}_{0})f\left(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{0},\rho\right)}$
$=\underbrace{\frac{f\left(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{a},\rho_{a}\right)}{f\left(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{0},\rho_{0}\right)}}_{A}\underbrace{\frac{f(\rho_{a}\mid\boldsymbol{\gamma}_{a})+\sum_{\rho\neq\rho_{a}}f(\rho\mid\boldsymbol{\gamma}_{a})\frac{f(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{a},\rho)}{f(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{a},\rho_{a})}}{f(\rho_{0}\mid\boldsymbol{\gamma}_{0})+\sum_{\rho\neq\rho_{0}}f(\rho\mid\boldsymbol{\gamma}_{0})\frac{f(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{0},\rho)}{f(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{0},\rho_{0})}}}_{B}$
where $\rho_{0}=\{\mathcal{I}_{1}\}$ and $\rho_{a}=\{\mathcal{I}_{1},\mathcal{I}_{2}\}$.
First focus on the $A$ term. Because of the orthogonality of the design matrix, from Som (2014) we know that
$\lim_{N\to\infty}\frac{f\left(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{a},\rho_{0}\right)}{f\left(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{0},\rho_{a}\right)}>0.$
Focus now on the $B$ term. From Theorem 4.3, the numerator is clearly strictly positive as long as the prior distribution $f(\rho\mid\boldsymbol{\gamma}_{a})$ puts non-zero probability on at least one refinement of of $\rho_{a}$. The prior distributions on partitions induced by the Dirichlet process obviously satisfies this requirement, as it places positive probability on every possible partition of $\mathcal{I}$. Similarly, for the denominator, note that the partitions over which we are summing are, by definition, all refinements of $\rho_{0}$. Hence, again from Theorem 4.3, we know that the limit of each of the terms in the sum is finite and strictly positive. Hence, $\lim_{N\to\infty}\sum_{\rho\neq\rho_{0}}f(\rho\mid\boldsymbol{\gamma}_{0})\frac{f(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{0},\rho)}{f(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{0},\rho_{0})}$ is also finite and therefore
$\lim_{N\to\infty}\frac{f(\rho_{a}\mid\boldsymbol{\gamma}_{a})+\sum_{\rho\neq\rho_{a}}f(\rho\mid\boldsymbol{\gamma}_{a})\frac{f(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{a},\rho)}{f(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{a},\rho_{a})}}{f(\rho_{0}\mid\boldsymbol{\gamma}_{0})+\sum_{\rho\neq\rho_{0}}f(\rho\mid\boldsymbol{\gamma}_{0})\frac{f(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{0},\rho)}{f(\boldsymbol{y}(N)\mid\boldsymbol{\gamma}_{0},\rho_{0})}}>0.$
This completes the proof.
Appendix F Proof of Theorem 4.5
Since
$Pr(\boldsymbol{\gamma}=\boldsymbol{\gamma}_{T}\mid\boldsymbol{y})=\frac{1}{1+\sum_{\boldsymbol{\gamma}\neq\boldsymbol{\gamma}_{T}}\frac{f(\boldsymbol{\gamma})}{f(\boldsymbol{\gamma}_{T})}\frac{f(\boldsymbol{y}\mid\boldsymbol{\gamma})}{f(\boldsymbol{y}\mid\boldsymbol{\gamma}_{T})}}$
and $f(\boldsymbol{\gamma}_{T})>0$, it is enough to show that
$\frac{f(\boldsymbol{y}\mid\boldsymbol{\gamma})}{f(\boldsymbol{y}\mid\boldsymbol{\gamma}_{T})}\xrightarrow[n\to\infty]{P}0.$
for all $\boldsymbol{\gamma}\neq\boldsymbol{\gamma}_{T}$. Now
$\frac{f(\boldsymbol{y}\mid\boldsymbol{\gamma})}{f(\boldsymbol{y}\mid\boldsymbol{\gamma}_{T})}$ $=\frac{\sum_{\rho}f(\boldsymbol{y}\mid\boldsymbol{\gamma},\rho)f(\rho)}{\sum_{\rho}f(\boldsymbol{y}\mid\boldsymbol{\gamma}_{T},\rho)f(\rho)}$
$=\underbrace{\frac{f(\boldsymbol{y}\mid\boldsymbol{\gamma},\rho_{0})}{f(\boldsymbol{y}\mid\boldsymbol{\gamma}_{T},\rho_{0})}}_{A}\underbrace{\left[\frac{f(\rho_{0})+\sum_{\rho\neq\rho_{0}}\frac{f(\boldsymbol{y}\mid\boldsymbol{\gamma},\rho)}{f(\boldsymbol{y}\mid\boldsymbol{\gamma},\rho_{0})}f(\rho)}{f(\rho_{0})+\sum_{\rho\neq\rho_{0}}\frac{f(\boldsymbol{y}\mid\boldsymbol{\gamma}_{T},\rho)}{f(\boldsymbol{y}\mid\boldsymbol{\gamma}_{T},\rho_{0})}f(\rho)}\right]}_{B}$
where $\rho_{0}=\{\{1,2,3,\ldots,p_{\boldsymbol{\gamma}}\}\}$, i.e., the partition that assigns all covariates to a single block.
Note that $A$ is the Bayes factor based on the standard $g$ prior. Since our hyperprior $p(g\mid a,b,\tau^{2})$ is a member of the Confluent Hypergeometric (CH) family of distributions and $\tau^{2}\sim\mathcal{O}(n)$, this Bayes factor is known to be consistent (e.g., see Li and Clyde, 2018). Hence,
$\frac{f(\boldsymbol{y}\mid\boldsymbol{\gamma},\rho_{0})}{f(\boldsymbol{y}\mid\boldsymbol{\gamma}_{T},\rho_{0})}$ $\xrightarrow[n\to\infty]{P}0$
for all $\boldsymbol{\gamma}\neq\boldsymbol{\gamma}_{T}$. On the other hand, $f(\boldsymbol{y}\mid\boldsymbol{\gamma},\rho)$ and $f(\boldsymbol{y}\mid\boldsymbol{\gamma},\rho_{0})$ share the same likelihood and differ only on their priors, which are both (approximately) unit information. Hence,
$\frac{f(\boldsymbol{y}\mid\boldsymbol{\gamma},\rho)}{f(\boldsymbol{y}\mid\boldsymbol{\gamma},\rho_{0})}\xrightarrow[n\to\infty]{P}c_{\boldsymbol{\gamma},\rho},$
for all $\boldsymbol{\gamma}$, where $0<c_{\boldsymbol{\gamma},\rho}<\infty$. Hence, $B$ converges to a finite constant, and the product of A and B converges to zero as desired.
Appendix G Details of the MCMC algorithm
[…]
  1. 1.
    We sample from the conditional posterior $f(\boldsymbol{\gamma},\tilde{\boldsymbol{g}},\boldsymbol{\xi}\mid\cdots)$ given by
    $f(\boldsymbol{\gamma},\tilde{\boldsymbol{g}},\boldsymbol{\xi}\mid\cdots)\propto f(\boldsymbol{y}\mid\boldsymbol{\gamma},\tilde{\boldsymbol{g}},\boldsymbol{\xi})f\left(\tilde{\boldsymbol{g}}\mid\boldsymbol{\gamma}\right)f(\boldsymbol{\xi}\mid\boldsymbol{\gamma},\alpha)f(\boldsymbol{\gamma}),$
    where $f(\boldsymbol{y}\mid\boldsymbol{\gamma},\tilde{g}_{1},\ldots,\tilde{g}_{K_{\boldsymbol{\gamma}}},\xi_{1},\ldots,\xi_{p_{\boldsymbol{\gamma}}})$ corresponds to (3.2) with $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}$, and $f(\boldsymbol{\gamma})$ is an appropriate prior on the space of models, e.g., a Beta-Binomial prior. We generate samples from the above distribution using a random walk Metropolis Hastings algorithm using a symmetric random walk proposal for $\boldsymbol{\gamma}$ similar to equation (46) of George and McCulloch (1997) as follows:
    • We define a probability vector $p_{1}=(0.7,0.3)$.
    • Each time, we decide on one of two types of moves according to the probability vector $p_{1}$.
      • If a move type 1 is selected, then the proposed new model $\boldsymbol{\gamma}^{(prop)}$ is generated by randomly flipping one component of $\boldsymbol{\gamma}$.
      • If a move type 2 is selected, the proposed model $\boldsymbol{\gamma}^{(prop)}$ is generated by removing one variable currently included in the model and replacing it with a variable currently excluded, leaving the dimensionality of the model unchanged. The variables to be added and removed are chosen uniformly at random within each set.
      If a new variable is included in the model $\{i:\boldsymbol{\gamma}=0,\boldsymbol{\gamma}^{(prop)}=1\}$, draw $\xi_{i}$ from the following distribution
      $\mathsf{Pr}(\xi_{i}=k)\propto\begin{cases}m_{\boldsymbol{\gamma},k}&\text{for }k=1,2,\dots,K_{\gamma},\\ \alpha&\text{for }k=K_{\gamma}+1.\end{cases}$
      If necessary, draw $\tilde{g}_{K_{\boldsymbol{\gamma}}+1}$ from the centering measure $f(\tilde{g}_{j}\mid\tau^{2},a,b)$.
      Similarly, if a variable is removed from the model, remove the corresponding $\xi_{i}$ and update the number of clusters and partition if a variable that was in a singleton cluster was removed. Update to get $\boldsymbol{\xi}^{(prop)},K_{\boldsymbol{\gamma}^{(prop)}},\rho_{\boldsymbol{\gamma}^{(prop)}}$ and $\tilde{\boldsymbol{g}}^{(prop)}$ accordingly. Then the proposed model is accepted with probability.
      $\min\left\{\frac{f(\boldsymbol{y}\mid\boldsymbol{\gamma}^{(prop)},\tilde{\boldsymbol{g}}^{(prop)},\boldsymbol{\xi}^{(prop)})p\left(\tilde{\boldsymbol{g}}^{(prop)}\mid\boldsymbol{\gamma}^{(prop)}\right)p(\boldsymbol{\xi}^{(prop)}\mid\boldsymbol{\gamma}^{(prop)},\alpha)p(\boldsymbol{\gamma}^{(prop)})}{f(\boldsymbol{y}\mid\boldsymbol{\gamma},\tilde{\boldsymbol{g}},\boldsymbol{\xi})p\left(\tilde{\boldsymbol{g}}\mid\boldsymbol{\gamma}\right)p(\boldsymbol{\xi}\mid\boldsymbol{\gamma},\alpha)p(\boldsymbol{\gamma})},1\right\}.$
  2. 2.
    Once the model is sampled, we can update $\beta_{0}$ and $\boldsymbol{\beta}_{\boldsymbol{\gamma}}$ by exploiting normal-normal conjugacy as follows:
    $\beta_{0}\mid\cdots$ $\sim\mathcal{N}(\bar{\boldsymbol{y}},\frac{\sigma^{2}}{n}),$
    $\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\cdots$ $\sim\mathcal{N}({\boldsymbol{m}}_{\boldsymbol{\gamma},\boldsymbol{\xi}},{\boldsymbol{V}}_{\boldsymbol{\gamma},\boldsymbol{\xi}})$
    where
    ${\boldsymbol{V}}_{\boldsymbol{\gamma},\boldsymbol{\xi}}=\sigma^{2}\left\{\frac{{\boldsymbol{G}_{\boldsymbol{\gamma}}}^{-1/2}\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}^{-1}{\boldsymbol{G}_{\boldsymbol{\gamma}}}^{-1/2}}{\tau^{2}}+\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1},\quad\quad{\boldsymbol{m}}_{\boldsymbol{\gamma},\boldsymbol{\xi}}=\frac{{\boldsymbol{V}}_{\boldsymbol{\gamma},\boldsymbol{\xi}}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{y}}{\sigma^{2}}.$
  3. 3.
    We can sample sample variance as
    $\sigma^{2}\mid\cdots\sim\text{Inverse-Gamma}\left(\frac{n-1}{2},\frac{\boldsymbol{y}^{T}(\boldsymbol{I}+\tau^{2}\boldsymbol{X}_{\boldsymbol{\gamma}}{\boldsymbol{G}_{\boldsymbol{\gamma}}}^{1/2}\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}{\boldsymbol{G}_{\boldsymbol{\gamma}}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T})^{-1}\boldsymbol{y}-n\bar{\boldsymbol{y}}^{2}}{2}\right)$
  4. 4.
    Conditional on the current model and the observational variance $\sigma^{2}$, sequentially sample $\xi_{i}$ for variables included in the model i.e. $\mathcal{I}=\{i:\gamma_{i}=1\}$ similar to Algorithm 8 of Neal (2000): Let $K_{\boldsymbol{\gamma}}^{-}$ be the number of distinct $\xi_{j}$ for $j\neq i$ and let $h=K_{\boldsymbol{\gamma}}^{-}+d$. We choose $d$ to be 20, by default. Label these $\xi_{j}$ with values $\{1,\dots,K_{\boldsymbol{\gamma}}^{-}\}$. If $\xi_{i}=\xi_{j}$, for some $j\neq i$, draw values independently from the base measure given by (3.3) for those $\tilde{g}_{k}$ for which $1\leq k\leq K_{\boldsymbol{\gamma}}^{-}$. If $\xi\neq\xi_{j}$ for all $j\neq i$, let $\xi_{i}$ have the label $K_{\boldsymbol{\gamma}}^{-}+1$ and draw independently from the base measure for those $\tilde{g}_{k}$ for which $K_{\boldsymbol{\gamma}}^{-}+1<k\leq h$. The, draw a new value of $\xi_{i}$ from $\{1,\dots,h\}$ with probabilities
    $\mathsf{Pr}(\xi_{i}=k\mid\cdot)\propto\begin{cases}m_{\boldsymbol{\gamma},k}^{-i}\phi\left(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\boldsymbol{0},\sigma^{2}\boldsymbol{G}_{\boldsymbol{\gamma},k}^{\star 1/2}\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma},k}^{\star 1/2}\right)&\text{ for }1\leq k\leq K_{\boldsymbol{\gamma}}^{-}\\ \frac{\alpha}{d}\phi\left(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\boldsymbol{0},\sigma^{2}\boldsymbol{G}_{\boldsymbol{\gamma},k}^{\star 1/2}\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma},k}^{\star 1/2}\right)&\text{ for }K_{\boldsymbol{\gamma}}^{-}+1<k\leq h\end{cases}$
    where $m_{\boldsymbol{\gamma},k}^{-i}$ is the number of $\xi_{j}$ for $j\neq i$ that are equal to k and $\boldsymbol{G}_{\boldsymbol{\gamma},k}^{\star}$ is same as $\boldsymbol{G}_{\boldsymbol{\gamma}}$ except for the fact that $g_{\xi_{i}}$ replaced by $g_{k}$.
  5. 5.
    Using equation (2) and (3) of Rodríguez (2013), the posterior distribution of the concentration parameter $\alpha$ can be written as
    $f(\alpha\mid\cdots)$ $\propto f(\boldsymbol{\xi}\mid\boldsymbol{\gamma},\alpha)f(\alpha\mid\boldsymbol{\gamma})$
    $\propto\frac{\Gamma(\alpha)}{\Gamma(\alpha+p_{\boldsymbol{\gamma}})}\alpha^{K_{\boldsymbol{\gamma}}}\prod_{k=1}^{K_{\boldsymbol{\gamma}}}\Gamma(m_{\boldsymbol{\gamma},k})\sqrt{\frac{1}{\alpha}\sum_{j=1}^{p_{\boldsymbol{\gamma}}-1}\frac{j}{(\alpha+j)^{2}}}$
    To sample from the above density, we employ a random walk Metropolis-Hasting algorithm with Gaussian proposals for $\log\alpha$; the default variance was the proposal was 0.05 but this needs to be tuned, depending on the dataset to achieve an average acceptance rate of 40-50%.
  6. 6.
    The conditional posterior distribution of $\tilde{g}_{k}$ for $k=1,\dots,K_{\boldsymbol{\gamma}}$ is given by
    $f(\tilde{g}_{k}\mid\cdot)$ $\propto\phi\left(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\boldsymbol{0},\sigma^{2}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right)f(\tilde{g}_{k}\mid\tau,a,b)$
    $=\phi\left(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\boldsymbol{0},\tau^{2}\sigma^{2}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\right)f(\tilde{g}_{k}\mid\tau=1,a,b)$
    We can simplify the above conditional posterior as
    $f(\tilde{g}_{k}\mid.)\propto{(\tilde{g}_{k})}^{b-\frac{m_{\boldsymbol{\gamma},k}}{2}}(1+\tilde{g}_{k})^{-a-b-2}\exp\left(-\frac{v_{k}}{\tilde{g}_{k}}-\frac{w_{k}}{\sqrt{\tilde{g}_{k}}}\right),$
    where
    $v_{k}=\frac{1}{2\sigma^{2}\tau^{2}}\sum_{\begin{subarray}{c}j\in S_{\boldsymbol{\gamma},k}\\ i\in S_{\boldsymbol{\gamma},k}\end{subarray}}\boldsymbol{\Sigma}_{\boldsymbol{\gamma},jj}^{-1}\beta_{\boldsymbol{\gamma},j}\beta_{\boldsymbol{\gamma},i}\quad\quad w_{k}=\frac{1}{\sigma^{2}\tau^{2}}\sum_{\begin{subarray}{c}j\in S_{\boldsymbol{\gamma},k}\\ i\notin S_{\boldsymbol{\gamma},k}\end{subarray}}\frac{\boldsymbol{\Sigma}_{\boldsymbol{\gamma},ji}^{-1}\beta_{\boldsymbol{\gamma},j}\beta_{\boldsymbol{\gamma},i}}{\sqrt{g_{\xi_{i}}}}.$
    Using the transformation $t_{k}=\frac{v_{k}}{\tilde{g}_{k}}$, we can re-parametrize this density as
    $f(t_{k}\mid.)\sim t_{k}^{a+\frac{m_{\boldsymbol{\gamma},k}}{2}}\left(1+\frac{t_{k}}{v_{k}}\right)^{-a-b-2}\exp\left(-t_{k}-\frac{w_{k}}{\sqrt{v_{k}}}\sqrt{t_{k}}\right).$
    Introduce the auxiliary variable $u_{k}$. Then, we can use slice sampling in conjunction with a modification of rejection sampler developed by Liu et al. (2012) to sample $t_{k}$ from a truncated extended gamma distribution to sample $t_{k}$ as follows
    $u_{k}\mid t_{k}$ $\sim\mathcal{U}\left(0,\left(\frac{v_{k}}{v_{k}+t_{k}}\right)^{a+b+2}\right),$
    $t_{k}\mid u_{k},.$ $\sim\text{Truncated-Extended-Gamma}\left(a+\frac{m_{\boldsymbol{\gamma},k}}{2}+1,\frac{w_{k}}{2\sqrt{v_{k}}},v_{k}(u_{k}^{\frac{-1}{a+b+2}}-1)\right),$
    where Truncated-Extended-Gamma distribution is given by
    $f(t\mid a,b,c)\propto t^{a-1}\exp(-t-2\sqrt{t}b)\mathbbm{1}_{\{0<t<c\}},\quad t>0,$
    for $a>0$ and $b\in\mathbb{R}$. Note that rejection sampler for untruncated extended gamma distributions developed by Liu et al. (2012) can be modified in a straightforward manner by using truncated proposals at the truncation level $c$. This can then be used to efficiently draw from truncated versions of extended Gamma distribution.
Appendix H Additional results for our second simulation study
In this section we supplement the results shown in Section 6.2 of the main paper by exploring the ability of the DP block $g$-prior to identify clusters of coefficients with similar shrinkage coefficients, as well as the estimation accuracy of the various procedures discussed in the paper.
(a) Dataset 31
(b) Dataset 63
Figure 9: Joint and marginal posterior distributions for $p_{\gamma}$ and $K_{\boldsymbol{\gamma}}$ in two representative examples of our simulated datasets in the $p=250$ and $\eta=0$ scenario.
First, to provide some intuition on the role of clustering in the performance of Dirichlet process mixtures of block-$g$ priors, we present in Figure 9, the joint and marginal posterior distributions for $p_{\gamma}$ and $K_{\boldsymbol{\gamma}}$ for the DP block $g$ ($\tau^{2}=n$) in two representative datasets. Note that, in both cases, the posterior distribution puts all of its mass in at least two clusters of coefficients, with a (marginal) mode of 3 and maximum of 9 in the Dataset 63, and mode of 6 and a maximum of 13 in Dataset 31. These results are in agreement with the theory developed in Theorem 4.3.
Next, Table 2 below is analogous to Table 1 of the main manuscript, but contains results for $\eta=0.5$ (a middle ground between the cases $\eta=0$ and $\eta=0.9$). These results are aligned with those presented in the main manuscript. Note that the difference between the GL $g$ and DP block prior when $\eta=0.5$ is very small, which is similar to what we obtained for $\eta=0$. This suggests that the differences between these two methods arise for relatively high correlations among the covariates.
Power (“large” coeffs) Power (“small” coeffs) Type I error (null coeffs)
$p=250$
$\eta=0.5$
g-prior ($\tau^{2}=$n) 0.979 0.793 0.005
Som et al. (K=2) 0.980 0.876 0.046
Som et al. (K=3) 0.990 1.000 1.000
GL-$g$ ($\tau^{2}=$1) 1.000 1.000 1.000
GL-$g$ ($\tau^{2}=$n) 0.987 0.869 0.060
GL-$g$ ($\tau^{2}\simHC$) 0.987 0.865 0.056
DP block-g ($\tau^{2}=$1) 0.990 0.903 0.246
DP block-g ($\tau^{2}=$n) 0.988 0.867 0.056
DP block-g ($\tau^{2}\simHC$) 0.988 0.876 0.102
ALasso 0.942 0.534 0.119
Horseshoe 0.985 0.844 0.026
HSM 0.982 0.813 0.006
$p=500$
$\eta=0.5$
g-prior ($\tau^{2}=$n) 0.968 0.721 0.000
Som et al. (K=2) 0.974 0.819 0.007
Som et al. (K=3) 0.981 0.901 1.000
GL-$g$ ($\tau^{2}=$1) 1.000 0.999 0.990
GL-$g$ ($\tau^{2}=$n) 0.979 0.812 0.008
GL-$g$ ($\tau^{2}\simHC$) 0.979 0.811 0.008
DP block-g ($\tau^{2}=$1) 0.980 0.826 0.025
DP block-g ($\tau^{2}=$n) 0.979 0.814 0.009
DP block-g ($\tau^{2}\simHC$) 0.980 0.817 0.011
ALasso 0.581 0.050 0.028
Horseshoe 0.974 0.781 0.011
HSM 0.974 0.765 0.001
$p=750$
$\eta=0.5$
g-prior ($\tau^{2}=$n) 0.942 0.479 0.000
Som et al. (K=2) 0.975 0.793 0.006
Som et al. (K=3) 0.980 0.844 0.502
GL-$g$ ($\tau^{2}=$1) 0.990 0.896 0.474
GL-$g$ ($\tau^{2}=$n) 0.980 0.783 0.004
GL-$g$ ($\tau^{2}\simHC$) 0.979 0.781 0.004
DP block-g ($\tau^{2}=$1) 0.980 0.796 0.008
DP block-g ($\tau^{2}=$n) 0.980 0.789 0.006
DP block-g ($\tau^{2}\simHC$) 0.980 0.788 0.007
ALasso 0.553 0.041 0.023
Horseshoe 0.973 0.725 0.005
HSM 0.976 0.739 0.000
Table 2: Estimates of power for “small” (generated from a $\mathcal{N}(0,1)$ distribution) and “large” (generated from a $\mathcal{N}(0,10)$ distribution) coefficients, and of type I error for null coefficients ($\beta=0$) in our second simulation study, for $\eta=0.5$. For the purpose of this table, coefficients are considered “significant” is their posterior inclusion probability is greater than 0.5.
Next, Figures 10 and 11 present the average mean squared error (AMSE) for the point estimators of the regression coefficients under various approaches (the posterior mean in the case of Bayesian procedures, and the argument of the penalized likelihood procedure for ALasso). As was the case for prediction MSEs in the main document, all results are shown relative to the AMSE associated with the standard $g$ prior. As before, ALasso is the worst performer, Horseshoe tends to perform poorly in “large pp” scenarios, and the remaining Bayesian procedures consistently outperform the $g$ prior (except, perhaps, for the null coefficients on some of the datasets). Perhaps surprisingly at first sight, Som et al. ($K=3$) outperforms the other procedures. While this result contrasts with what happened when carrying out variable selection (where Som et al. ($K=3$) was one of the worst performers), it is not surprising. The overfitting of the shrinkage factors for null coefficients that led to issues in model selection becomes an advantage when it comes to the accuracy of point estimators.
Finally, we study the sensitivity of procedures based on global-local $g$-priors and Dirichlet process mixtures of block $g$-priors to the hyperpriors on the concentration parameter $\alpha$ and on the vector of inclusion indicators $\boldsymbol{\gamma}$. To this effect, Table 3 shows the power and type I error associated with DP block-$g$ ($\tau^{2}=n$) and DP block-$g$ ($\tau^{2}\sim\mathsf{HC}$) where $\alpha$ is given an Exponential distribution with mean 1, as well as results for GL-$g$ ($\tau^{2}=n$) and DP block-$g$ ($\tau^{2}=n$) under a uniform prior on model space. Note that the prior on $\alpha$ seems to have no effect on the power or type I errors. On the other hand, the use of a uniform prior on model space seems to lead to slightly higher power for detecting smaller coefficients when $p\ll n$ and slightly lower power when $p\geq n$. On average, this results in lower $F_{1}$ scores under a uniform prior on model space.
Figure 10: Relative mean squared error of the coefficients for $\eta=0$, broken down by coefficient block. Results are shown in the log scale because of the poor performance of ALasso.
Figure 11: Relative mean squared error of the coefficients for $\eta=0.9$, broken down by coefficient block. Results are shown in the log scale because of the poor performance of ALasso.
Power (large coeffs) Power (small coeffs) Type I error (null coeffs) Power (large coeffs) Power (small coeffs) Type I error (null coeffs)
$p=250$
$\eta=0$ $\eta=0.9$
DP block-g ($\tau^{2}=$n) Gamma 0.990 0.900 0.043 0.974 0.755 0.111
DP block-g ($\tau^{2}/n\simHC$) Gamma 0.990 0.902 0.045 0.976 0.772 0.159
GL-$g$ ($\tau^{2}=$n) Uniform 0.987 0.867 0.008 0.962 0.626 0.018
DP block-g ($\tau^{2}=$n) Uniform 0.987 0.867 0.007 0.964 0.638 0.022
$p=500$
$\eta=0$ $\eta=0.9$
DP block-g ($\tau^{2}=$n) Gamma 0.984 0.867 0.006 0.958 0.611 0.018
DP block-g ($\tau^{2}/n\simHC$) Gamma 0.984 0.868 0.005 0.958 0.614 0.019
GL-$g$ ($\tau^{2}=$n) Uniform 0.985 0.874 0.010 0.959 0.635 0.024
DP block-g ($\tau^{2}=$n) Uniform 0.986 0.879 0.018 0.963 0.661 0.047
$p=750$
$\eta=0$ $\eta=0.9$
DP block-g ($\tau^{2}=$n) Gamma 0.985 0.854 0.003 0.950 0.534 0.009
DP block-g ($\tau^{2}/n\simHC$) Gamma 0.985 0.853 0.003 0.949 0.537 0.010
GL-$g$ ($\tau^{2}=$n) Uniform 0.987 0.875 0.017 0.962 0.631 0.058
DP block-g ($\tau^{2}=$n) Uniform 0.988 0.882 0.085 0.965 0.654 0.132
Table 3: Sensitivity analyses. Estimates of power for small (generated from a $\mathcal{N}(0,1)$ distribution) and large (generated from a $\mathcal{N}(0,10)$ distribution) coefficients, and of type I error for null coefficients ($\beta=0$) in our second simulation study. For the purpose of this table, coefficients are considered “significant” is their posterior inclusion probability is greater than 0.5.